### v1

# An Industry Interview Study of Software Signing for Supply Chain Security

Abstract
Many software products are composed by the recursive integration of components from other teams or external parties. Each additional link in a software product’s supply chain increases the risk of the injection of malicious behavior. To improve supply chain provenance, many cybersecurity frameworks, standards, and regulations recommend the use of software signing. However, recent surveys and measurement studies have found that the adoption rate and quality of software signatures are low. These findings raise questions about the practical application of software signing, the human factors influencing its adoption, and the challenges faced during its implementation. We lack in-depth industry perspectives on the challenges and practices of software signing.
To understand software signing in practice, we interviewed 18 high-ranking industry practitioners across 13 organizations. We provide possible impacts of experienced software supply chain failures, security standards, and regulations on software signing adoption. We also study the challenges that affect an effective software signing implementation. To summarize our findings: (1) We present a refined model of the software supply chain factory model highlighting practitioner’s signing practices; (2) We highlight the different challenges – Technical, Organizational, and Human — that hamper software signing implementation; (3) We report that expert subjects disagree on the importance of signing; (4) We describe how failure incidents and industry standards affect the adoption of software signing and other security techniques. Our findings contribute to the understanding of software supply chain security by highlighting the impact of human and organizational factors on Software Supply Chain risks and providing nuanced insights for effectively implementing Software Supply Chain security controls — towards Software signing in practice.
1 Introduction
Like all engineered products, software can be built more quickly when components are reused from other sources [1, 2]. This integration accelerates production but requires trusting external parties [3]. This trust increases when the provenance of components is known [4]. The strongest guarantee of provenance is a cryptographic signature created using a secret known only to the vendor. This technique, known as software signing, is underutilized by many software producers [5, 6]. However, as a result of software supply chain attacks such as SolarWinds, signing has become integral to many government, academic, and industry proposals to secure software supply chains [7, 8, 9].
Although software signing has desirable properties, it must be adopted carefully to bring benefits. At present, measurement studies have shown that missing or erroneous signatures are common across the software supply chain [5, 10, 11, 6, 12]. When signing is omitted or done unsuccessfully, researchers have demonstrated exploits [8, 13, 14, 15]. Existing security frameworks recommend software signing, but the ambiguous guidelines across these frameworks lead to a lack of a detailed implementation model [16, 17]. In order to adopt software signing, organizations would benefit from a detailed understanding of industry practices and challenges. The academic and grey literature does not offer detailed industry perspectives on the experiences that engineers have when adopting software signing.
Our work represents the first qualitative study of software signing in industry practice. We interviewed 18 industry security experts and decision-makers across 13 organizations. Our interviews focused on four aspects: the software signing practices employed in industry; the importance accorded to software signing in practice; the influence of industry standards and failure events; and the organizational, technical, and other challenges that hamper the effective implementation of software signing in practice.
We present a comprehensive summary of our findings. We refined the software supply chain factory model to show the roles of signing. While artifacts are frequently signed, signature verification is notably low. Our study reveals that the most common challenges in practice are technical, organizational, and human factors. These include key management and infrastructure, compatibility, a lack of signature verification, expertise in setting up signing tools, and developer attitudes. Our research indicates that software signing is typically used in one of three ways to mitigate risks: establishing provenance, serving as an auxiliary security technique, or fulfilling requirements akin to ticking a checkbox. While software supply chain failures, regulations, and other industry standards do influence the adoption of software signing and other security techniques, our participants have identified several issues associated with these regulations and standards.
Our contributions are:
  1. 1.
    We refined the software supply chain factory model to reflect diverse industry practices and the forms that signing may take therein (Figures 4 and III).
  2. 2.
    We identified common technical, human, and organizational challenges in adopting signing (Table IV).
  3. 3.
    We report on the relative importance of signing in software supply chain security strategies (section 5.3).
  4. 4.
    We discuss the effect that failures, standards, and regulations have on software signing practices (section 5.4).
Significance: Software signing is a widely recommended technique, and has become mandatory for US federal software vendors. Many organizations are therefore seeking to integrate software signing into their engineering processes. This study is the first qualitative work on software signing in industry. We provide insights on software signing in industry: adoption practices, challenges, and human and organizational factors to consider. Such knowledge will guide organizations in their software signing journey, and contribute to the software engineering community’s strategies for enhancing software supply chain security.
2 Background & Related Works
Here, we cover software supply chains (section 2.1), cybersecurity considerations (section 2.2), and empirical studies (section 2.3).
2.1 Software Supply Chains
Software production is often depicted using variations on the early Software Factory Model [18, 19] or the more recent Software Supply Chain Factory Model (SSCFM) [20, 21], depicted in Figure 1. The SSCFM approximates software product from the perspective of one software engineering organization or team. The SSCFM model shows that a software product integrates first-party and third-party components, that these components are packaged for downstream use, and that the software product is at risk of malicious code injection from internal or external parties.
Figure 1: The Software Factory Model [20].
The study of software supply chains has been driven by engineering organizations’ greater reliance on external components. Many works have reported on this in various software contexts. For example, a 2024 Synopsys report stated that, in 1,067 projects across 17 industry sectors, $\sim$96integrated open-source components, and $\sim$77of total code is from open source [22]. Even in government- and safety-critical systems, the use of third-party open-source components has moved from anathema to normal [23, 24]. Analyses of open-source software package registries have shown large increases year-over-year in the number of available components, the number of actively-used components, and the number of transitive dependencies [25, 26, 27, 28].
When an external component contains a vulnerability, that vulnerability may be propagated to its dependents. The increased reliance on external components in software production has amplified the potential for exploitation of these vulnerabilities, and many attack scenarios have been proposed [29, 30, 31, 32, 33]. For example, Lauinger et al. [34] report that $\sim$37of the most popular Internet websites rely on at least one library with a publicly acknowledged vulnerability. Stock et al. showed that among 1,273 real-world client-side XSS vulnerabilities, 273 cases occurred exclusively within third-party code [35]. Upstream in the supply chain, Zimmerman et al.’s analysis of npm indicated that up to 40% of all packages depended on code with at least one publicly known vulnerability [26]. Given the prevalence of such inherited vulnerabilities, Zahan et al. [36] and similar studies have summarized attack vectors and potential weak links in software supply chains [37, 38, 39, 40, 41, 42, 43, 44].
2.2 Securing Software Supply Chains
2.2.1 Methods and Practices
Numerous methods and tools have emerged to improve software supply chain security. Okafor et al. [45] characterize them as promoting three distinct dimensions: transparency (knowing all components in the chain), separation (isolating components), and validity (ensuring integrity of components). For example, transparency is increased by adoption of Software Bills of Materials (SBOMs) [46, 47]. Separation is promoted by techniques such as sandboxing [48, 49, 50, 51, 52] and zero-trust architectures [53, 54]. Validity is improved by reproducible builds [55, 56], software signatures [5], and attestations about artifact metadata such as authorship [57].
Security standards, frameworks, and regulations have been established to complement and direct these security efforts. Notable examples include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [20], the Cloud Native Computing Foundation’s (CNCF) Software Supply Chain Best Practices [58], Microsoft’s Supply Chain Integrity Model (SCIM) [59], and the US National Institute of Standards and Technology’s (NIST) Secure Software Development Framework (NIST-SSDF) [60] and NIST 800-204D [61], which were developed in conjunction with US Executive Order 14028 (EO-14028) [62]. At present, national regulations are inadequate to force compliance with secure supply chain practices [63].
2.2.2 Software Signing in Detail
Software signing is a formally guaranteed method of establishing the authorship of a software component [64, 65, 66, 14]. Signing ensures the provenance of components through public-key cryptography (Figure 2), improving validity (provenance) in supply chains [45]. Many standards and government regulations recommend signing as a base Software Supply Chain security technique [20, 7, 67, 59, 9]. For example, in The Linux Foundation’s SLSA framework, software signing is needed for Build Security Level 3 [20].
Figure 2: Typical workflow for software signing and verifying signatures. The component author packages (A) and signs (B) their software. The signed package (C) and public key (D) are published. To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
2.3 Empirical Work on SW Supply Chain Security
The theoretical properties of techniques to improve software supply chain security are well known, whether via transparency, separation, or validity. However, adopting them in practice remains a challenge. Empirical/in vivo studies on software supply chain security have examined adoption practices and challenges to identify gaps in theoretical guarantees or usability. Some studies examine the general activities of component management, e.g., studies on dependency selection [68] and maintenance [69], and examinations of security challenges and incident handling procedures [70]. Others have examined specific techniques for supply chain security. For example, with respect to the transparency property, researchers have studied the practices of SBOM adoption [71, 72], and reported gaps in tooling and interoperability as well as incomplete adoption and maintenance across the software development lifecycle.
With respect to software signing, researchers have examined gaps in usability and adoption. Older implementations of signing were complex to use [73, 74], inspiring newer approaches with simplified tooling and support for key management [75, 76, 77]. About 25% of publicly-reported attacks from 2010-2021 exploited the missing or erroneous use of signing [15]. Gokkaya et al. showed that the top signing-related attack vectors were self-signed code, broken signature systems, and stolen certificates [42]. Several reports show that signing adoption rates for open-source components are low. Zahan et al. found that $\leq$0.5of packages in the npm and PyPI registries had signed releases [6]. Schorlemmer et al. corroborated this proportion, further noting that signing adoption is correlated with registry policies and that the correctness or quality of signatures depends on usable tooling [5]. Ladisa et al.’s survey shed light on these measurements, finding that software signing is valued but is considered moderately costly to set up [44, 78].
3 Knowledge Gaps & Research Questions
In our analysis of the literature discussed in section 2.3, we observed that our empirical understanding of software signing challenges and adoption is primarily derived from open-source software components. The available data on industry practices are surveys [44, 78]. We lack in-depth qualitative descriptions of industry experiences, creating a gap in our understanding of software supply chain security. We do not know which signing strategies are effective, the challenges associated with implementing these methods, the roles played by suggested methods and tools, and the factors influencing their adoption. To fill this gap, we ask four research questions:
  • RQ1:
    How and why is software signing implemented by software teams?
  • RQ2:
    What are the challenges that affect the implementation and use of software signing?
  • RQ3:
    What is the perceived importance of Software Signing in mitigating risks?
  • RQ4:
    How does software supply chain failures and standards influence the adoption of software signing?
4 Methodology
We conducted semi-structured interviews with subjects who were cybersecurity practitioners. We depict our methodology in Figure 3. In this section we discuss our research design rationale (section 4.1), protocol development (section 4.2), participant recruitment (section 4.3, data analysis (section 4.4), and study limitations (section 4.5).
Figure 3: Study methodology. Several sources (left) informed our protocol topics (center). We applied thematic and framework analyses (right).
4.1 Approach — Research Design Rationale
Our questions require insight into practitioner perspectives. We considered two factors in selecting a methodology. First, since the prior work in this space is limited, there was relatively little basis on which to structure a closed-ended survey. Second, we expected recruitment to be a problem, since the target population has high security expertise and a broad organizational view, and thus are likely to be high-ranking (and busy) people. We therefore opted for a semi-structured interview [79], which would permit us to capture emergent results [80] (factor 1), and to leverage our professional relationships to improve response rates (factor 2).
The resulting data is expected to be exploratory, and rooted in practitioners’ perspectives and experiences [79]. The phenomenon of interest is software signing as a security method, which is a contextual decision based on an organization’s distinct policies, security expectations, perceived threats, and customer requirements [81]. We therefore adopt an interpretivist perspective for our research methodology design and analysis. An interpretivist approach treats each situation as unique [82], focusing on in-depth variables and factors related to each context [83].
4.2 Protocol Design and Development
To create the initial interview protocol, we studied three kinds of resources. First, to ensure a common language and a comprehensive set of topics, we studied software supply chain security frameworks such as SLSA, CNCF, SCIM, and NIST-SSDF (section 2.2). These frameworks use variants of the Software Supply Chain Factory Model (SSCFM, cf. Figure 1). We therefore used the SSCFM as our reference framework for structuring the C Sections of our interview protocol. This choice allowed us to investigate signing practices at various stages of the SDLC while using a model familiar to our subjects. Second, we examined grey literature sources to gain insights into some practical uses of software signing. Our primary reference was the official Sigstore blog as of Jan. 31, 2023. This resource has eight user reports about adopting software signing in various industry contexts, from which we derived questions about the contextual and organizational factors that subjects consider. Third, we reviewed adjacent literature [68, 69, 84] to understand effective questions related to software supply chain security, while avoiding duplication.
After protocol development, we refined the protocol through practice interviews (2) and pilot interviews (2). For practice, the lead author interviewed two of the secondary authors. These practice sessions led to modifications in sections B and C of our interview protocol (see Appendix Appendix C). After recruiting subjects, we treated the first two subjects as pilot interviews. During these pilots, we gathered additional insights into how subjects perceived software supply chain attacks and reorganized the protocol, particularly focusing on questions related to the perceived sources of greatest risk (section B). Since few questions were changed or added, we included the data from the pilot interviews in our study where feasible. The final change occurred after the seventh interview. The author team discussed the results so far and identified one emergent topic: how practitioners’ organizations integrate feedback into their SSC security processes. One question was added for use in the remaining interviews.
Our interview protocol is summarized in Table I, with a full version in Appendix Appendix B.
TABLE I: Summary of interview protocol. See Appendix Appendix B for full protocol. Topic D was not analyzed in this study.
Topic (# Questions) Sample Questions
A. Demographics (4) What is your role in your team?
B. Perceived software supply chain risks (4) Can you describe any specific software supply chain attacks (e.g., incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
C. Mitigating risks with signing (8) How does the team use software signing to protect its source code (what parts of the process is signing required)?
D. Signing tool adoption (5) What factors did the team consider before adopting [TOOL/METHOD] over others?
4.3 Data Collection
Our Institutional Review Board (IRB) approved this study.
TABLE II: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives. “Technical leader” — senior, lead, partner, and principal engineers. “Engineer” and “Manager” — junior staff.
ID Role Experience Product Area
S1 Research leader 5 years Digital technology
S2 Senior mgmt. 15 years SSC security
S3 Senior mgmt. 13 years SSC security
S4 Technical leader 20 years Cloud, Dev tools
S5 Engineer 2 years Cloud security
S6 Technical leader 27 years Digital technology
S7 Manager 6 years Cloud security
S8 Technical leader 8 years Internet services
S9 Engineer 2.5 years SSC security
S10 Engineer 13 years Dev tools
S11 Technical leader 16 years Social technology
S12 Technical leader 4 years Cloud, OSS security
S13 Senior mgmt. 16 years Telecommunications
S14 Research leader 13 years Cloud security
S15 Senior mgmt. 15 years Aerospace security
S16 Technical leader 26 years Digital technology
S17 Senior mgmt. 15 years SSC security
S18 Manager 11 years Internet services
Target Population: To address our research questions effectively, the target population was industry practitioners with expertise in software signing. An ideal subject would have in-depth knowledge and experience related to software supply chain practices within their organizations. Such experts would have high credibility, insider perspectives, and nuanced interpretations.
Recruitment: We employed a non-probability-based Purposive and Snowball population sampling approach to recruit such subjects [82]. Our initial subject sample pool was members from the Cloud Native Computing Foundation’s Kubecon 2023 conference organizing committee, many of whom also contribute to software supply chain security projects hosted by The Linux Foundation. This pool contained many security experts, although we acknowledge that the connection to specific signing projects may bias their responses to this genre of signing approach. Subsequently, we expanded our subject pool through recommendations from this initial group of subjects and other industry contacts. We offered each participant a \$100 gift card as an incentive.
In total, we contacted 30 candidates, of whom 18 agreed to participate (60% response rate). These subjects came from 13 different organizations. Their demographic information is summarized in Table II. Subjects were affiliated with companies with security products or were part of their company’s security team. For subject anonymity, we do not distinguish those from CNCF or from snowballing.
Interviews: Interviews were conducted by the lead author via Zoom. The median interview duration was 50 minutes.
4.4 Data Analysis
Interview recordings were transcribed by www.rev.com (human transcription). We anonymized identifying information. Our analysis then proceeded in several stages.
(Stage 1) We began with data familiarization [85]. Three analysts (A1–A3) undertook an inductive reading, memoing, and coding process for all transcripts.
(Stage 2) Subsequently, following O’Connor & Joffe’s recommendation [86] to initially independently code 10-25% of the data, we randomly selected approximately 33% (6 transcripts) for multiple memoing. Two analysts (A1, A2) conducted this multiple memoing of data in three rounds, collaborating to discuss emerging code categorizations after each round. This process yielded 506 coded memos, which were reviewed by Analysts A1 and A3 over four one-hour meetings to create a codebook comprising 36 code categories, excluding demographics.
(Stage 3) To ensure the reliability of our coding, analysts A1 and A3 randomly selected and coded a transcript. We used the Percentage agreement metric as recommended by Feng [87] for easy coding tasks with nominal data. Since our codebook was collectively defined by both analysts A1 and A3, we assumed this coding task should not be too difficult — we achieved an 89% agreement rate, confirming this assumption. Any discrepancies in coding were resolved through discussion and code redefinition.
(Stage 4) Next, a second set of 6 transcripts was selected and coded by Analyst A1 using an initial codebook, adding 6 minor categories, each with less than 6 codes. This updated codebook was reviewed by Analyst A3.
(Stage 5) The same process was applied to the final 6 transcripts, yielding 4 more minor categories with fewer than 4 codes each. Out of the 10 new categories, only 2 provided new insights into our data, as determined through multiple review rounds by Analysts A1 and A3.
(Thematic and Framework Analyses) From this process we induced themes following Braune & Clark’s method [88]. These themes helped us answer RQs 2-4 directly. To answer RQ1, we also applied a framework analysis [89], mapping the themes to the software supply chain factory model as a reference framework to analyze participants’ responses and further categorize the themes.
Through our coding phase’s thematic analysis, we identified eight themes. Four of these themes were focused on the practical implementation of Signing, while the remaining four addressed various aspects of software signing: the challenges encountered, its perceived importance, the impact of security failures on its adoption, and the influence of security standards and regulations on its adoption.
To determine whether additional aspects of the phenomenon might be expected to emerge if we continued to sample, we measured saturation in the 18 interview transcripts. Following Guest et al.’s [90] recommendation, we tracked the cumulative appearance of new codes in each interview. Each interview had a median of 33 unique codes, also depicted in  Figure 5. Saturation was achieved after interview 11 (details in Appendix Appendix D).
4.5 Study Limitations
Positionality Statement: We acknowledge the role that our backgrounds may have played in this study [91]. The author team are cybersecurity researchers with expertise in software signing, so our expertise both enabled and influenced the protocol development and analysis. Beyond this, one of the authors is a contributor to a popular signing tool. To reduce bias in our study, that author had no involvement in the development of the research protocol or analysis. However, that author reviewed the results and framing of the study, to increase its validity for the software signing community.
Other Threats to Validity: This work is qualitative, and bears the limitations of any qualitative study. First, the cost of data collection reduces the sample size, threatening generalizability. Our subject recruitment may also have introduced biases, e.g., towards CNCF/The Linux Foundation-oriented frameworks and tools. As mitigations, we note that our codebook saturated at the 11th interview, with few/no new codes observed in subjects 12-18 (Appendix D); and our sample size is typical for similar studies [71, 72]. Second, qualitative data implies subjective analysis. To improve the reliability of our results, we followed an iterative process to develop the codebook and used multiple raters with a high measure of agreement.
Specific to the topic of software signing, the frameworks for software supply chain security and the available automation for implementing signing are evolving. Our results should be understood as a snapshot of current practices in industry. They will grow out of date as signing standards, regulations, and tooling continues to evolve.
5 Results
5.1 RQ1: Signing Implementations in Practice
Finding 1. We refine the software supply chain factory model from Figure 1, as shown in Figure 4 and Table III. Our refined model shows candidate points for the establishment of provenance. We analyze our subjects’ responses using this refined model, identifying points on this model where subjects’ teams or organizations employ software signing.
Figure 4: Our refined software supply chain factory model (compare to Figure 1), showing different points where provenance and integrity of software could be established (PI, PE, PB, PD, PS, PA) and Verified (VI, VE, VB, VD, VS, VA) using software signatures. Abbreviations are defined in Table III.
TABLE III: Different Signing use Points as noted in  Figure 4 and subjects who employ software signing at these points. We mark subjects who consider them optional for reasons of demand by customers, nature of products, etc.
Point Description Subjects
SOURCE CODE
PI Internal code contributors sign their commits S2-S11, S14-S17 (S4, S8, S14 – Optional)
PE External code contributors sign their commits S7, S10 (S11, S14 — Optional)
VI Verification of signatures from Internal contributors S7, S14, S17
VE Verification of signatures from external contributors
PSC Signing after code reviews/audits S2-S11, S14-S17 (S4, S8, S14 — Optional)
BUILD
VSC Verification of source code signatures before build S2, S3, S7, S14
PB Signing of build output S1-S5, S7, S8, S11-S15, S17 (S8 —Optional)
DEPLOYMENT/PACKAGE/ARTIFACT
VB Verification of build signature before Deployment or final artifact S7, S14
PA Signing of Final Software product All S1–S18
THIRD-PARTY DEPENDENCY
PD Signing after verification and certification of dependencies S6, S7, S9, S16, S17, S18
VD Verification of Signatures from external dependencies S9, S11, S17 (S9, S11, S17 — Optional)
VA Verification by final customers Difficult to determine
Using our initial analysis plus the recommendations and philosophy of several frameworks, which recommend a continuous verification of integrity and trust in code alone (not humans) [20, 92], we present a refined form of the software supply chain factory model as shown in Figure 4. Our model comprehensively identifies points in the software delivery process where the security property of provenance is expected to be established or verified. We compare this with subjects’ reported usage of other security methods for these factory model components, and second with framework (and standards) prescriptions.
5.1.1 Source code
First, we ascertained whether our subjects used any form of external code contributors in their software creation process. Out of 10 subjects who reported a form of collaboration with open-source or other external contributors, only 5 subjects used external code contributions in the production code of projects they owned (both commercial and OSS). We then proceeded to determine if any form of signing is used to establish provenance for points PE, PI, VE, VI, and PSC.
External contributors sign commits (PE): Signing is mostly required by teams who use external code contributors. Out of 5 subjects whose teams used external code contributions in their production code, 4 of these subjects require that commits be signed. However, two of those do not require signing for open-source projects or noncritical projects. Only 2 of our subjects unconditionally require signing from their external contributors. As S2 put it, “I got something signed from somebody I don’t know…So we do the code review, and then we have a sign-off by somebody on our team that signs off from that code”.
Internal contributors sign commits (PI): Signing is almost always required by teams for their internal code contributors. 14 of 18 subjects report that their team required some sort of commit signing from internal contributors. Subjects who list this as an optional requirement (3) cite similar reasons as in PE above.
Verification of signatures from contributors (VE&VI): While signing is relatively required, it is not verified as a check for security and provenance of code contributions from contributors. Techniques like code reviews (and security audits), security/vulnerability scans, tests, etc, are the most used security methods for assessing the security of code contributions. S10 opined to this, “I think everybody on the team does sign their commits, but we don’t really verify it…If somebody stopped signing their commits, I don’t think much would really happen”
Code is signed after reviews/audit (PSC): Signing here is common similar to PI. Subjects who report that signing is required in PI, also report that all activities related to code reviews, or audits are (cryptographically) signed off on.
5.1.2 Build
Next, we extract from our subjects’ responses how signing is used in the build process of software products. For most of the subjects, the build and build signing (for subjects who do so) are automated. Our model expects that the source code to be built should be verified (VSC), and then when the build is successful we expect another round of provenance certification (PB).
Verification of source code signatures before build (VB): Signatures from source codes are rarely verified before the commencement of the build process. Two subjects mentioned some sort of rigorous verification at this point. Alternate means of security at this stage are similar to the security techniques reported for the build process.
Signing of build output (PB): 13 subjects reported the use of signing at this stage of the software delivery process. 10 subjects reported signing is used alongside attestations, SBOMS, and other forms of metadata.
5.1.3 Package/Artifact
Next, we analyze the use of signing at the final stage of deploying software products.
Verification of build signature before deployment of final artifact (VB): Similar to the other verification steps, signatures from the build phase are rarely verified before the deployment. Only two subjects mentioned verification at this point. For example, S7 states, “We sign our code and then we also verify that it was built in our specific build system”.
Other subjects mentioned that the security of the entire build process comprises techniques like an automated build pipeline with Workflow definitions, use of build policies, and Reproducible builds. These are complemented by code scanning tools as in the case of the source code stage.
Signing of Final Software product (PA): All 18 subjects said that signing is done at the deployment stage to establish provenance for the organization on each of their products.
5.1.4 Third-Party Dependencies
For third-party dependencies in the Software Supply Chain, we evaluated how software signing is used as a factor influencing dependency selection, and how software signing is used to verify the authenticity of dependencies in practice.
Presence of a Signature as a factor for Dependency Selection: For most of our subjects, signatures on a third-party dependency do not affect their choice in selecting such a dependency. Only 3 of our subjects like to see this, treating signatures as a signal that the owner is security conscious. Other security factors that influence the selection of third-party dependencies as mentioned are: the use of the OpenSSF’s scorecard tool, the dependency’s known vulnerabilities, level of activity, licensing, and age.
Verification of Signatures from external dependencies (VD): Signatures on third-party dependencies are rarely verified in practice. Only three subjects reported some form of verifying signatures from third-party dependencies. As S11 said, “If it is open source, then probably no…But if it is commercial off the shelf, we want to make sure those are signed.” However, most of our subjects reported several methods for assessing the authenticity of a third-party dependency. These methods include code and vulnerability scanning (e.g., Coverity), source rebuilding, internal mirroring of dependencies, and reliance on ecosystem internal checks (e.g., Go package manager’s checks). Other subjects perform internal reviews of dependencies, while some have a centralized dependency selection and management process, where dependencies are hosted and maintained by the organization. A few subjects noted that their organization discourages or forbids third-party dependencies.
Signing after verification and certification of dependencies(PD): A few subjects reported that third-party dependencies are thoroughly reviewed using one of the aforementioned techniques before being signed off for use by a team member. Six of our subjects indicated that this process is followed in their software production workflow. For example, S17 described their team’s process like this: “We have a signed report that’s signed by us, and we trust us…that says, ‘Hey, I scanned this third-party dependency, did all these things. I don’t have a signature for that, but I ran all these extra checks and then signed those checks. So now we have a record”’. In contrast to that attestation approach, S18’s team ingests third-party dependencies and signs the internal versions directly.
5.2 RQ2: Challenges affecting the implementation of software signing in practice
Finding 2. Subjects identified several common themes as challenges faced in practice as obstacles to effectively implementing software signing. We categorize these broadly; 1. Technical Challenges: These include – key Management and Infrastructure issues, compatibility issues with existing systems and different architectures, and lack of signature verification 2. Organizational and human Challenges: These include – expertise in setting up Signing implementation, operationalization of the signing procedure, developer attitude, and Resources to set up the signing process
Our next research objective was to understand the various challenges practitioners face while implementing software signing in practice. We grouped subjects’ responses under three main themes for challenges faced by practitioners in practice; Technical, Organizational, and Human factors. Table IV gives a summary. Due to space limitations, we focus on only a subset of the individual challenges.
TABLE IV: Challenges to software signing Implementation in practice. We have categorized related challenges highlighted by subjects. Two uncommon Technical challenges are described in Appendix E.
Observed Challenges #Subjects(#Orgs) Proposed Solutions
TECHNICAL
Key management 10 (9) Use of Keyless Signing (e.g., Sigstore)
Compatibility issues 6 (6)
Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
Ease of use/Usability 4 (4) Usable signing tools (e.g., Sigstore), Documentation
No unifying Standard 2 (2)
Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
ORGANIZATIONAL
Operationalization of the Signing process 4 (4) Automating Signing
Resource to Set up Signing 3 (3)
Creating Effective Signing Policy 2 (2) Regular Process Feedback mechanisms
No Management incentive to Sign 2 (2)
Bureaucracy 1 (1)
HUMAN
Expertise in setting setup and use of signing 5 (4)
Developer Attitude to Signing 3 (3) Automating Signing
Lack of Demand from Customers 1 (1)
5.2.1 Technical Challenges
Key management: Key management Issues were the most reported issues. Issues under this category vary from key management, key distribution, key generation, and public key infrastructure (PKI) issues.
Key management issues are the highest categories recorded. S13 articulates this, especially concerning key distribution: “a lot of the ways in which I think previous teams’ software signing hasn’t been useful is…key management…in terms of how the private key is managed but also in how the public key is made available.”
Some subjects discussed recent signing technologies that implement keyless signing (e.g., Sigstore) as a possible solution. S18 said: “Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have these big servers to store all these signatures”. S17 also noted, “[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada. So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better.”.
Compatibility issues: Compatibility issues arise during the integration of software signing systems and tools with other tools within the software engineering process, including other security tools and the artifacts to be signed.
S2, S7, S17, and S18 highlight issues surrounding the compatibility of newer signing systems with existing software systems in the software engineering process. As S18 put it, “There are teams who’ve been doing this for quite a while…Selling that is a big challenge, right? Because we need to prove that it is much more secure than what has been there in place.”
S8, whose organization’s software product is built for different platforms and hardware architectures, noted compatibility issues across various architectures: “Every platform evolved separately…macOS…Windows…The requirements are changing and evolving independently”.
Compatibility with other security tools can pose challenges too, as noted by S5: “The vulnerability scan report [is for]…code…CVE published…[but] signatures, I don’t think it will pick those things up.”
Ease of use/Usability: Usability of the signing tool or implementation was a common challenge hampering effective use in practice. This was often expressed as developer experience, e.g., S17 said: “For a developer, make sure that they can sign the commits easily”. This point was also raised at a team level by S13: “I think that awareness of the value of signing and sophistication around how signing is performed has increased significantly, especially in recent years. But it has also resulted in a lot of smaller teams deciding that this is not a manageable thing.”
With respect to documentation, S11 said: “I think there are multiple challenges there. When you build some signing infrastructure, you want to make it a self-serve so the teams can onboard themselves, how well you have your self-serve playbooks written down, documentation on using those. I think documentation can be improved.”
Lack of Verification of Signatures: Lack of verification of signatures is also quoted by several subjects as a main challenge. This arises from the factors like signing serves as a “checkmark” security tool (S4: “most people who are strictly signing all commits don’t have a meaningful process for verifying those signatures or for enforcing any policy around them. They just do it because”), and lack of customer interest (S14: “We have tried to … get potential customers and customers interested…but as far as I know, none of our customers…verify our signatures on images”).
Verification Of Artifact before Signing: Verification generally was highlighted as a problem, especially verifying inputs to the build process of the software delivery process. The problem noted here is that the content of software artifacts is not readily verified before signing: as S16 put it, “if you start with the component, the component gets put into a subassembly. The subassembly gets put into a package. The package can get put into a broader package or whatever. I think the people are signing. But do you know that the sum of all the pieces are accurate? … what are you really signing, and at what level?”. S11 further noted that different teams could sign arbitrary artifacts due to shared signing infrastructure.
5.2.2 Organizational Challenges
Operationalization of the Signing process: The implementations of software signing are typically neither uniform nor centralized across different organizations. Various teams are largely responsible for crafting their own solutions. As S6 said, “I can’t be more specific, but today, I’ve talked to three different teams with three different processes”. S15 suggested that automation could improve uniformity: “We try to break down the barriers with really good examples, by having common configuration of CI through similar GitHub workflows, and stuff like that. So it’s easier for people to use.”
Resource to Set up Signing: The issue of organizational resources to set up efficient signing tooling across an organization was also another cause for concern for some of our subjects. Subjects discussed this in terms of finances, time, and engineering effort. S14 said, “We need to do a lot of signing…hundreds of container images [daily]…It was a pretty significant engineering effort to build a pipeline that can handle all those signatures.”
Creating Effective Signing Policy: Policies surrounding the implementation of software signing by teams are generally not considered highly effective. These policies include elements such as role-based access, key usage, and key issuance. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “… the other thing is policies, creating a policy that is, for one, it’s hard to do. And then the other part is creating an effective policy and knowing that it’s effective”
A suggested solution to address this issue is the use of regular feedback mechanisms to assess the security policies and processes of different teams within the organization. As S11 said, “We want to be more proactive…internal audits, and an assessment methodology like the NIST, I think CISF.”
No Management incentive to Sign: A lack of management incentives was also a challenge. S11 said, “The investment, the time that the engineer is putting into…signing…before a release, does it get the right attention from the senior management? Do they view it as a time well spent for the product?”
5.2.3 Human Challenges
Expertise: Some subjects noted that setting up signing tools depends on the engineer’s expertise. Unfortunately, subjects felt that most engineers lack the necessary experience and skills for this task. As S9 said, “there’s a lift there as far as…expertise”.
Developer Attitude to Signing. Software developers’ attitudes toward software signing pose challenges to the signing process. Disinterest and resistance to adopting new tools are the primary attitude-related obstacles cited. S10 observed: “Most developers don’t really understand this stuff and don’t really want to think about it…[they are] happy if they…see a big green tick[mark]”. S6 added: “They refuse to get off their old tools”
Lack of Demand from Customers: Most customer-driven organizations report that software signing is challenging to implement because customers rarely request this feature in their requirements. S14 put it like this: “Sign[ing] our images was there from day…we haven’t found many customers that are too excited about the signing aspect”.
5.3 RQ3: Perceived importance of Software Signing in mitigating risks
Finding 3. Subjects disagreed on this point. Some view signing as crucial, while others treat it as a secondary consideration in their overall strategy.
Here, we aimed to understand how important practitioners perceived software signing to be in their software delivery security. Opinions vary on the significance of signing in the software delivery process.
5.3.1 As Crucial to Provenance and Integrity
Many of our subjects considered provenance an integral part of security, and felt that software signing is (or will be) a guarantee of that. S4 said, “I think signing is massively important. Signing for open source is probably the single most important thing for software supply chain security”.
5.3.2 As a Secondary Security technique
Some subjects believe that software signing is a supplementary method used to ensure the integrity of primary security techniques. They often deemed it important only in conjunction with methods like Attestations and other Manifest forms such as the SBOM. As S2 said, “I think the most important thing is accurate metadata collection and centralization. And even if that data’s not signed, we can still get a lot of usefulness out of it. Signatures, in my opinion, they offer an additional layer of security on top of what we should be doing, which is metadata collection”. S17 similarly views signing as a final check on top of other security practices: “Signing was essentially the way of saying, ‘I have checked, we did all the [security practices]’.”.
5.4 RQ4: Impact of Supply Chain Security Awareness on Software Signing Practices
Finding 4. 1. Software supply chain failures usually require immediate fixes to the malfunctioning or compromised components. Large-scale failures typically trigger a comprehensive reassessment, leading to changes in the security strategy of a team or organization. 2. Security regulations, frameworks, and standards drive the adoption of signing and other security techniques, but some subjects criticized these obligations.
In this section, we explore how experience with software supply chain security influences a company’s decision to adopt and implement software signing. We asked subjects about their personal and organizational experiences with software supply chain failures (attacks, vulnerabilities, and other incidents) and how these experiences affected their team’s or organization’s software supply chain security decisions, particularly regarding software signing. Additionally, we queried subjects on the impact of software supply chain security frameworks, standards, and regulations on their use of software signing and other security techniques.
5.4.1 How experiences with SSC attacks influence the use of software signing
TABLE V: Experienced software supply chain failures and type of Security change triggered.
Type # Subjects (# Orgs) # Direct fix # $\Delta$Signing # $\Delta$General Security process
Attack 4 (4) 4 2
Vulnerability 9 (6) 9 1 2
Other Incidents 5 (5) 5 2
Software supply chain security failures typically have direct fixes, except in cases of severe attacks. While only 4 of our subjects indicated experiencing software supply chain attacks, only one of these attacks (where the subject was hired to help fix the aftermath) was highlighted as major, resulting in collateral damage and leading to a complete overhaul of the software and security processes. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S4: “I would say that at this point it’s probably difficult to find any process at that organization that wasn’t inspected and potentially changed as a result of that. Every aspect of how everything was built there was changed … There was no choice in that situation.”
Some subjects report that software supply chain security failures led to a broader change in their team’s or organization’s security posture. For example, S18 describes their team’s response to the Log4j vulnerability, which was present in their system. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… it was a reality check to know that even companies like ours … can be attacked through open source libraries, … Asset inventory is the biggest thing that is lacking where you don’t know what is currently running in your environment. And if it is running, you don’t know how old that is, or what are the dependencies which is currently running in production. So that pushed us towards making sure that we have a continuous inventory of container images, which are currently running, for example, for Kubernetes, and then there are initiatives like GUAC.”
Generally, most software supply chain reported by our subjects were not severe, and as such most of them received direct fixes. This could be swapping out a defaulting tool or dependency, re-releasing an affected release, publishing security advisories, etc. We highlight some of those responses. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S5: “So what happens is that, if the current version where that security risk is present, if that is also publicly available, we have to create a security advisory to let everyone know that if you are working with this version of our software, you should upgrade to the next version which contains the fix. And we let them know what the fix is and if they are affected by the fix.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “to my knowledge it did not cause a major restructuring or change that is … Yeah, I would say that they probably patched any of the security issues that they had.”
The only instance where a signing-related fix directly resulted from a software supply chain security failure was when the subject team’s product was a safety-critical application. This team added intoto’s signed attestation capability. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S2: “we started reaching for the In-toto project, and we saw that it allowed you to decouple the security metadata collection and validation. You can abstract that away from the actual build process itself. So that decoupling allowed us to solve this problem of how do we know that this opaque process for the software, that opaque certification, how do we know that was followed without actually knowing the details of what that process is?”
5.4.2 Influence of security regulations, frameworks, and standards
TABLE VI: Regulatory and Standard Influences on Security Technique Adoption: Analysis of Subjects’ Responses. 4 Subjects say they are not influenced by either.
ANY INFLUENCE?
#Subjects (Yes) #Subjects (None)
Regulation(EO-14028) 11 7
Other Standards 8 10
The adoption of software signing and other security techniques is heavily influenced by regulations and other standards. Specifically, we asked our subjects, “Are these strategies (software signing implementation & other complementary security methods(earlier listed by subjects)) influenced by regulations and standards?”. While answering generally, most of our subjects agreed that regulations, particularly the E0-14028, played a key role in the way their team and organization have implemented software supply chain security. Although software signing is not explicitly mentioned, Some subjects did note that their use of SBOMs was directly a result of regulations or standards. We only had 1 subject who noted that the standards influenced their use of signing in certifying attestations(SBOMs).
S1 highlights this influence of the executive order: “Yeah, I would say certainly the collection of SBOMs and doing that more broadly has been a response to things like the executive order, at least for our team specifically. I would say, yeah, the executive order played just a strong role in what we are starting to do now. Whether or not some teams are now trying to be preventative, that’s a good question.” S7 also adds to S1’s point: “we keep the executive orders in mind when we build our product…”
Some of our subjects (3) only relied on standards (not on regulations) for security recommendations. For example: {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… as an organization, currently we don’t have any government contracts. And I think the executive order is mostly targeted towards companies, organizations who are having contracts with the government and they need to produce certain SBOMs and things like those. So our push towards supply chain or better security was not influenced by the executive order.”
S18 also adds: “But I think in terms of where we are at, I think in terms of frameworks, as you mentioned, we do look at some of these frameworks like SLSA,… There are quite a few …, and each of them target different parts of the supply chain. But as an organization, it’s really difficult to keep track and follow each of these standards of frameworks, to be honest, because we have limited resources, we can’t just keep on handling those. So we are looking at SLSA, or at least trying to get closer to where we should be for the SLSA level. Again, it’s not written in stone that we need to have level one, level two, level three, things like that, but that’s something which we strive to achieve sooner.”
S14 adds to this line of thought: “The recent interests from the US federal government and software supply chain security, SBOMs, and the executive order from 2021, it’s hard to detect very much influence on our security because of those … I think when observers of software security read that executive order and then commentary in the aftermath, they often think that SBOMs have become mandatory…”
5.4.3 Criticism of Regulations and Standards
Also, our subjects identified several issues with the current set of software supply chain security standards and regulations that hinder their influence on the adoption of software signing and other security tools. We highlight four kinds of issues.
(a) Regulations don’t specify standard data exchange format for security tools: Criticism of standards and regulations here lies in their inadequacies to specify or promote standard data interchange formats. This is highlighted by the lack of data formats for certain current security techniques such as Software Bill of Materials (SBOMs), other manifest-type techniques, signing tools, etc. The absence of standards for these techniques leads to numerous undesired practices, such as the use of vulnerability scan reports as an input for SBOM data. This ultimately leads to compatibility problems among different security techniques.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S16: “…one of the gaps with the executive order is, yeah, you give an SBOM now. But what are they going to do with it? Nobody knows yet. And how do you compare it? Because even right now, going back to the tools, there are different versions of SPDX. There are different versions of Cyclone DX. They’re the two big ones. This goes back to tools. There are a bunch of tools out there that will convert say 2.0 back to three and you come up with different answers using different tools.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “SBOMs are a good example of something that has a standard, but all of the implementations end up being different enough that they’re very difficult to translate from one format to the other or even get consistency within the same format type.”
(b) Regulation and standards are just checkbox rules:
Here, subjects argue that standards and regulations function more as compliance checks rather than as security enhancers. They point out that these standards and regulations lack security enhancement features, such as implementation models. Instead, they primarily promote compliance-type features, such as guidelines and best practices lists. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S17: “So there is currently a disconnect between regulations and security. So compliance and regulations keep people out of jail, keep people from getting fined, [while] security stops you from getting hacked. And the problem there is sometimes those two things can conflict. And compliance is there to make sure that everybody is following reasonable sets of rules and security is there to actually implement that and create the software and the processes that hit that … Regulations help provide the requirement. They don’t necessarily make security better on their own. And there’s always going to be a balance because what compliance tells you is compliance helps you with making sure that everybody’s following the standard … the only way for me to be compliant to this is to ignore the actual problem and just check the box.”
(c) Difficult to implement in the SDLC: A group of participants highlighted the challenges encountered when attempting to integrate the recommendations of security standards into their software engineering processes. These challenges primarily arise from the difficulties in training engineers on the application of these standards, and the deficiencies in the technical implementation of the standards themselves. S12 makes this point clear in their response; {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S12: “… just having the executive order was just another way to explain it to the higher-ups… I don’t think that the executive order in particular affected the business … There’s kind of milestones. You don’t need to know exactly what the ins and outs are technically of achieving them … But it’s a lot of work getting that in place and upskilling all of your existing developers, et cetera. So yeah, anything that requires going to all 4,000, 10,000 of your developers in some of these massive organizations and making sure every single one of them does something different, it is a lot of work.”
(d) Standards are not formalized: Criticism regarding the formalization of standards, as expressed by some participants, reflects the lack of consensus surrounding these standards. Participants argue that since these standards are not necessarily created by a majority of industry consortia, they should not be referred to as ‘standards’. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S13: “the term standards is imprecise. So some people think of standards as being very formalized but some people would say any documented practice is a standard. So there’s the SLSA project in it, and one might argue that that is a standard or one might argue that it’s not a standard, but it is a documented set of recommendations. And if people are implementing those, you could argue that they’re implementing a standard.”
6 Discussion
6.1 Recommendations from this study
Improving Signature & Artifacts Verification in Signing workflows.   Our refined software supply chain factory model analysis has highlighted a significant gap in the current practice of signing - signature verification. This issue was underscored by several subjects in our study.
In certain automated workflows, the act of signing is often treated as a mere formality, serving as a “checkbox” that, once done, allows engineers to carry out their usual tasks (e.g., creating pull requests or merging code into the main branch). Thus, this checkbox provides very little security properties, and requires effort to provide meaningful security gains.
Further, while artifact verification, in this context, refers to conducting a thorough check of the integrity of the software artifact by means of identifying the signer. However, the use of signed attestations (i.e., are statements about a particular action or property in the supply chain) can further verify the authenticity and integrity of the process and will become crucial to further improve the security posture of organizations.
Future works in this area should focus on proposing improved tooling around existing or new software signing tools to include these capabilities in their design. This enhancement would not only improve transparency but also reduce the costs and expertise requirements for setting up such infrastructure.
By addressing these issues, we can ensure that signing is not just a procedural element, and instead plays a crucial role in maintaining the integrity and authenticity of the software development process.
Addressing Incentive Challenges in Software Signing.  One of the most frequently reported challenges in the realm of software signing is the lack of proper incentives for its adoption. These incentives can range from support from management, the ease of using signing tools, as well as comprehensive education on the importance and application of signing.
While some of these issues, such as the need for educating practitioners on the significance and application of signing, can be addressed with relative ease, others necessitate more elaborate planning. For instance, fostering management support might require demonstrating the tangible benefits of software signing in terms of enhanced security and reduced risk of software tampering.
Moreover, though the user-friendlness of signing software has been evidenced by various “why Johnny can’t” papers, more research is needed in lowering the adoption bar for software signing systems.
Regardless of the complexity of these challenges, investing in these areas is paramount for enhancing an organization’s security posture. By fostering a culture that values and understands the importance of software signing, organizations can ensure the integrity of their software and protect against potential security threats.
Future research could focus on developing strategies to address these incentive-related challenges and exploring their effectiveness in different organizational contexts. This could provide valuable insights for organizations striving to improve their software signing practices.
Criticism of Current Standards and Regulations.  A significant finding of our study is the criticism directed towards current standards and regulations, particularly concerning their creation process and the difficulty of their implementation. These issues could be mitigated by ensuring that these standards and regulations are informed by a broader section of the industry consortium. This would involve gathering input from a diverse range of practitioners, including those from different sectors, roles, and levels of expertise.
While frameworks like SLSA already involve collaboration among multiple organizations, we advocate for a more inclusive effort. This would target a larger group of practitioners, ensuring that the standards and regulations are not only comprehensive but also practical and implementable across various contexts.
In the future, research could focus on exploring effective strategies for fostering such inclusive collaborations and assessing their impact on the quality and applicability of the resulting standards and regulations. By addressing these issues, we can work towards standards and regulations that are not only robust and comprehensive but also widely accepted and easily implementable. This would significantly enhance the security and integrity of software signing practices across the industry.
6.2 Contrast to prior work
Comparisons with Previous Quantitative Studies on Software Signing.  We compare our findings with Schorlemmer et al.’s [5] quantitative study on software signing in open-source ecosystems. Their results indicate that major software supply chain failures (e.g., SolarWinds, Codecov attacks) and dedicated tooling had no significant impact on the quantity of signatures. In contrast, our findings for the non-open source industry show the opposite. This difference is expected due to the substantial incentives, such as commercial and reputation losses, that drive greater adoption in the non-open source industry. Additionally, our study supports this finding, as some subjects noted that contributions to the open-source ecosystem often involve relaxed security requirements.
Distinctions of Refined Model Compared to Existing Frameworks.  While all software supply chain security frameworks agree on the defined threat model, our refinement of this model differs from existing frameworks in several ways.
  1. 1.
    SLSA: The SLSA framework primarily focuses on securing the build process within the software engineering lifecycle, offering recommendations as a series of incremental security hardening levels. In contrast, our refinement addresses the entire software supply chain, not just the build step.
  2. 2.
    SCIM and S2C2F: This framework provides threat-reduction recommendations without specific implementation details. Our approach, however, emphasizes the continuous establishment and verification of provenance across the software supply chain.
  3. 3.
    Hammi et al. [93]: While their work proposes a blockchain-based cryptographic tool, we do not advocate for a specific tooling solution. Instead, our model is designed for industry and software teams to create an implementation workflow that is independent of specific tools. Additionally, our findings are supported by insights from industry practitioners, enhancing the practical applicability of our model.
7 Conclusion
Signature-based supply chain provenance is a topic of increasing importance, driven by cyberattacks, executive orders, and new industry standards. Although the literature contains many studies of signing as a technology, there have been few studies of the use of signing in engineering contexts. Those that exist have focused on open-source, or on the mechanics of signing. In this paper, we present the first qualitative examination of signing in engineering organizations’ security strategies. We interviewed 18 high-ranking practitioners across 13 organizations, including C-suite staff, principal and senior engineers, managers, and consultants. Our four main findings are that (1) Software products are being signed, but the signatures are not yet being validated; (2) The factors that hamper signing include infrastructure and compatibility issues, lack of expertise, developer and organizational attitudes, and lack of resources; (3) Signing is considered an important guarantee of provenance, but it is not the main technique these organizations consider in their cybersecurity postures; and (4) Organizational behaviors are being influenced by major failures (e.g., SolarWinds) and by regulations and industry standards, but some of our subjects expressed concern about the development process and quality of those recommendations. Our results indicate many opportunities for technical and operational advances in the use of software signing to improve software supply chain provenance.
8 Replicability
For anonymity reasons, we do not plan to share the raw transcripts of our interviews. However, our interview protocol, codebook, etc. are available in the Appendices.
9 Acknowledgments
We thank the study participants for contributing their time to this work. Taylor Schorlemmer and Wenxin Jiang assisted in typesetting. We acknowledge support from Google, Cisco, and NSF #2229703.
Appendix A Outline of Appendices
The appendix contains the following material:
  • Appendix B: The interview protocol.
  • Appendix C: The codebooks used in our analysis, with illustrative quotes mapped to each code.
  • Appendix D: Analysis of saturation in our data.
  • Appendix E: Uncommon challenges, omitted from the main paper due to space limitations.
Appendix B Interview Protocol
Table I gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
We include Section D of the protocol for completeness, although it was not analyzed in this study.
TABLE VII: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions.
Themes Sub-themes Questions
A: Demographic A-1: What best describes your role in your team? (Security engineer, Infrastructure, software engineer, etc)
A-2: What is your seniority level? How many years of experience?
A-3: What is the team size?
A-4: $\dagger$ What are the team’s major software products/artifacts?
A-5: What type of Organization/company?
B: Software Supply Chain
Failure Experienced
B-1: Describe briefly the team’s process from project conception to product release and maintenance – This is to understand the unique context of each practitioner’s software production case.
B-2: $\star$ What do you consider a Software Supply chain attack/incident to be?
B-3: Can you describe any specific software supply chain risks (or incidences with third-party dependencies, code contributors, open source, etc.) that your team has encountered during your software development process? How were these addressed? (Alternatively – How did this affect the decision to implement software signing?
[…]
B-5: What is your team’s greatest source of software supply chain security risk? Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from  LABEL:sec:_B-sec-experienced): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
Third-party dependencies C-2: What are the team’s peculiar selection strategies for Third-party dependencies? How does the presence of a signature influence this decision? How is the authenticity of this signature verified? Any other methods/practices to ensure the trustworthiness of the third-party dependencies before integrating them into your projects? Possible Follow-up – Can you describe these methods/practices?
C-3: What influences the team to discontinue the use of a package? How does the Signature on the dependency influence this decision?
C-4: How does the team manage its third-party dependencies’ security? How do you maintain awareness of potential vulnerabilities or threats related to third-party components?
First-party source code C-5: Is software signing a requirement for team developers (insider threats) and open-source contributors (if any)? how does the team use software signing to protect its source code (what parts of the process is signing required e.g. Commit signing)?
Build Process C-6: (How) Does the team utilize signing in its build process?
Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism?
D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
D-2: What factors did the team consider before adopting this tooling over others?
D-3: What was the team’s previous signing practice before the introduction of current tooling?
D-4: How does your team implement this tooling (which components of tool does the team use)?
D-5: Have you encountered any challenge(s) using this tool of choice? How have you coped with this limitation(s)? Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
D-6: Have you/your team considered switching this tool for another? AWhat other tools have been considered or are currently being considered?
Appendix C Codebooks
Table VIII describes our codebook, with code, definition, and example quote tagged with that code.
TABLE VIII: Excerpts from the final codebook we used for our analysis. Sample codes under each theme give examples. Note that the theme, “Others and Specific tooling” is not discussed in the paper.
Code Definition Sample Quote
BUILD
Build Infrastructure of team Nature of Subject’s build process or build Infrastructure. Example- when subject mentions that their build or CICD is automated S7: "Our build system is typically GitHub actions."
SOURCE CODE
Signing Implementation in Source Code (External Contributors) Subject’s manner of implementing Software signing in first-party source code produced by external contributors like 3rd party vendors, example, commits by contributors are signed S2: "I don’t know who these people are … I got something signed from somebody I don’t know … So what we do though is that we do the code review, and then we have a sign-off by somebody on our team that signs off from that code"
DEPLOYMENT/PACKAGE/ARTIFACT
Signing Implementation in Package/Artifacts Subject’s manner of implementing Software signing in Deployment of Product or packaging of created artifact (Release) S10: "git commit signing on the code and then signatures on the final artifact. "
THIRD-PARTY DEPENDENCY
Signing Implementation in Third Party Dependencies and Tools Contains code category related to 1. Influence of Signing in Selection of third party dependencies and 2. Influence of Signing in discontinuation of dependency use S13: "So most software package ecosystems that I consume packages from do not support signing consistently enough to put much stock in that. "
EFFECT OF SSC FAILURE ON ADOPTION
Experienced Software Supply Chain Attack Subject fell victim to a malicious orchestrated software supply chain attack. For example, an organization falling for phishing S4: "The build system itself was compromised by a state actor…somebody compromised the system and lurked for a long time and built very bespoke malware, specifically tailored to Orion, to mimic Orion’s network traffic, all of that other kind of stuff…NSA and GHGQ both said that this was Russian foreign intelligence that did this to SolarWinds."
SPECIFIC TOOLING
Sigstore Criticisms Criticism and weakness associated with implementing SigStore for signing S3: "we do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey…removing that intermediary…to do identity-based signing directly against the OIDC provider."
PERCEIVED IMPORTANCE OF SIGNING
Provenance and Integrity Subjects mention that their current or previous teams use software signing in a major role for establishing provenance and integrity or other similar roles. S4: "A few years from now, everybody expects provenance…they don’t install anything if they can’t prove where it came from…I think signing is massively important. Signing for open source is probably the single most important thing for software supply chain security"
SSC STANDARDS AND REGULATION INFLUENCE
Influence of Regulations and Standards on SSC Security Strategies How regulations and standards influence their SSC security strategies on signing, or use of other security methods S8: "My external looking in perspective of things, is that there’s a mix of anticipation of regulation, as well as reaction to regulation. For the most part, there’s a lot of trying to predict what is going to come, what the government is seeking for as far as tech regulation, and trying to get ahead of that curve."
CHALLENGES
Challenges experienced in the use of Software Signing These are responses from subjects when asked about challenges to them implementing Software Signing generally and not just criticism of a signing implementation like Sigstore/GPG/Notary etc. For example when the subject talks about Key management, Bureaucracy, motivation S7: "I’m not confident that those customers actually ever verify those signatures."
OTHERS
Choice of tool - General Reasons What influences choice of software supply chain security tools except signing tools S17: "A lot of the tools are not particularly easy to use "
Appendix D Saturation Analysis
Figure 5 shows the saturation curve associated with our interviews.
Figure 5: Saturation curve. Each interview covered many topics in detail (orange line, blue line). However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
Appendix E Uncommon Challenges
E.0.1 Managing Signed Artifacts
Managing signed software artifacts post-signing should follow good practices, but this is currently a challenge. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S9: “Also managing all the artifacts that are signed … It’s great that you sign things, but what do you do with it afterwards? I think that’s kind of something that we’re facing right now, especially when it comes to container management. Great, we signed containers, but now what? … Okay, great, now what? All those downstream things that we’re trying to figure out.”
E.0.2 Identity Management/Authentication
Managing and authenticating the identities of team members authorized to sign at various stages of the software delivery process poses challenges in some instances.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “in general one of the challenges that I hear from other security architects and security leads, and program managers, and people at that level, is key management and identity management as being a barrier for signing and security and authentication in general.”
Possible solutions suggested and used by some participants include the use of short-lived or constant key rotation. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S7: “We really rely on short-lived tokens for authentication. I think that is one way that we really try to restrict access to our system and make sure that there’s no long-lived secrets, there’s no long-lived access tokens anywhere in our system. For any signing that we do, I think those keys are rotated quite frequently. So, that’s another thing we do to prevent any attack should key information be leaked. I’m trying to think what else.”

### v2

@@
 Abstract
-Many software products are composed by the recursive integration of components from other teams or external parties.
+Many software products are composed of components integrated from other teams or external parties.
 Each additional link in a software product’s supply chain increases the risk of the injection of malicious behavior.
@@
 However, recent surveys and measurement studies have found that the adoption rate and quality of software signatures are low.
-These findings raise questions about the practical application of software signing, the human factors influencing its adoption, and the challenges faced during its implementation.
 We lack in-depth industry perspectives on the challenges and practices of software signing.
-To understand software signing in practice, we interviewed 18 high-ranking industry practitioners across 13 organizations.
-We provide possible impacts of experienced software supply chain failures, security standards, and regulations on software signing adoption.
-We also study the challenges that affect an effective software signing implementation.
-To summarize our findings: (1) We present a refined model of the software supply chain factory model highlighting practitioner’s signing practices; (2) We highlight the different challenges – Technical, Organizational, and Human — that hamper software signing implementation; (3) We report that expert subjects disagree on the importance of signing; (4) We describe how failure incidents and industry standards affect the adoption of software signing and other security techniques.
-Our findings contribute to the understanding of software supply chain security by highlighting the impact of human and organizational factors on Software Supply Chain risks and providing nuanced insights for effectively implementing Software Supply Chain security controls — towards Software signing in practice.
+To understand software signing in practice, we interviewed 18 experienced security practitioners across 13 organizations.
+We study the challenges that affect the effective implementation of software signing in practice.
+We also provide possible impacts of experienced software supply chain failures, security standards, and regulations on software signing adoption.
+To summarize our findings: (1) We present a refined model of the software supply chain factory model highlighting practitioner’s signing practices; (2) We highlight the different challenges–technical, organizational, and human–that hamper software signing implementation; (3) We report that experts disagree on the importance of signing; and (4) We describe how internal and external events affect the adoption of software signing.
+Our work describes the considerations for adopting software signing as one aspect of the broader goal of improved software supply chain security.
 1 Introduction
-Like all engineered products, software can be built more quickly when components are reused from other sources [1, 2].
-This integration accelerates production but requires trusting external parties [3].
-This trust increases when the provenance of components is known [4].
-The strongest guarantee of provenance is a cryptographic signature created using a secret known only to the vendor.
-This technique, known as software signing, is underutilized by many software producers [5, 6].
-However, as a result of software supply chain attacks such as SolarWinds, signing has become integral to many government, academic, and industry proposals to secure software supply chains [7, 8, 9].
-Although software signing has desirable properties, it must be adopted carefully to bring benefits.
-At present, measurement studies have shown that missing or erroneous signatures are common across the software supply chain [5, 10, 11, 6, 12].
-When signing is omitted or done unsuccessfully, researchers have demonstrated exploits [8, 13, 14, 15].
-Existing security frameworks recommend software signing, but the ambiguous guidelines across these frameworks lead to a lack of a detailed implementation model [16, 17].
+Like all engineered products, software is built more quickly when components are reused from other sources [75, 40].
+Although this integration accelerates production, it requires trusting external parties [5].
+This trust increases when the provenance of components is known [68], such as via software signing.
+Although historically underutilized [64, 97], due to attacks such as SolarWinds [31], signing has become part of many government, academic, and industry proposals to secure software supply chains [80, 19, 17].
+Although software signing has desirable effects, it must be adopted carefully to bring benefits.
+At present, measurement studies show that missing or erroneous signatures are common across the software supply chain [64, 53, 42, 97, 93].
+Existing security frameworks recommend software signing, but the general guidelines in these frameworks omit a detailed implementation model [61, 100].
 In order to adopt software signing, organizations would benefit from a detailed understanding of industry practices and challenges.
-The academic and grey literature does not offer detailed industry perspectives on the experiences that engineers have when adopting software signing.
-Our work represents the first qualitative study of software signing in industry practice.
-We interviewed 18 industry security experts and decision-makers across 13 organizations.
-Our interviews focused on four aspects: the software signing practices employed in industry; the importance accorded to software signing in practice; the influence of industry standards and failure events; and the organizational, technical, and other challenges that hamper the effective implementation of software signing in practice.
-We present a comprehensive summary of our findings.
-We refined the software supply chain factory model to show the roles of signing.
-While artifacts are frequently signed, signature verification is notably low.
-Our study reveals that the most common challenges in practice are technical, organizational, and human factors.
-These include key management and infrastructure, compatibility, a lack of signature verification, expertise in setting up signing tools, and developer attitudes.
-Our research indicates that software signing is typically used in one of three ways to mitigate risks: establishing provenance, serving as an auxiliary security technique, or fulfilling requirements akin to ticking a checkbox.
-While software supply chain failures, regulations, and other industry standards do influence the adoption of software signing and other security techniques, our participants have identified several issues associated with these regulations and standards.
-Our contributions are:
+This view is currently lacking in the academic and grey literature.
+Our work represents the first interview study that describes software signing practices and rationales in industry.
+This method gives a deep look at a topic, exposing many directions for further study and implications for practice.
+In this work, We interviewed 18 experienced security practitioners across 13 organizations.
+Our interviews focused on four aspects: the software signing practices employed in industry; the challenges that hamper the effective implementation of software signing in practice; the importance accorded to software signing in practice; and the influence of failure events and industry standards.
+Our results provide a detailed understanding of industry practices and challenges in using software signing to promote software supply chain security.
+By integrating our subjects’ descriptions, we introduce a refinement of the Software Supply Chain Factory Model that indicates the possible and actual roles of software signing.
+Our results reveal three kinds of challenges in adopting signing: technical, organizational, and human aspects.
+Additionally, our findings indicate that practitioners hold varying views on the importance of signing for supply chain security.
+Lastly, we found that security events — both major failures such as SolarWinds, and new standards and regulations — appear to have a minimal effect on the adoption of software signing.
+In summary, we make the following contributions:
   1. 1.
-    We refined the software supply chain factory model to reflect diverse industry practices and the forms that signing may take therein (Figures 4 and III).
+    We refined the software supply chain factory model to reflect diverse industry practices and the forms that signing may take therein (section 5.1).
   2. 2.
-    We identified common technical, human, and organizational challenges in adopting signing (Table IV).
+    We identified common technical, human, and organizational challenges in adopting signing (section 5.2).
   3. 3.
@@
     We discuss the effect that failures, standards, and regulations have on software signing practices (section 5.4).
-Significance: Software signing is a widely recommended technique, and has become mandatory for US federal software vendors.
-Many organizations are therefore seeking to integrate software signing into their engineering processes.
-This study is the first qualitative work on software signing in industry.
-We provide insights on software signing in industry: adoption practices, challenges, and human and organizational factors to consider.
-Such knowledge will guide organizations in their software signing journey, and contribute to the software engineering community’s strategies for enhancing software supply chain security.
 2 Background & Related Works
-Here, we cover software supply chains (section 2.1), cybersecurity considerations (section 2.2), and empirical studies (section 2.3).
+Here, we cover software supply chains (section 2.1), cybersecurity considerations (section 2.2), and relevant empirical studies (section 2.3).
 2.1 Software Supply Chains
 Software production is often depicted using variations on the early Software Factory Model [47, 14] or the more recent Software Supply Chain Factory Model (SSCFM) [79, 1], depicted in Figure 1.
-The SSCFM approximates software product from the perspective of one software engineering organization or team.
-The SSCFM model shows that a software product integrates first-party and third-party components, that these components are packaged for downstream use, and that the software product is at risk of malicious code injection from internal or external parties.
-Figure 1: The Software Factory Model [20].
+The SSCFM approximates software product from the perspective of one software engineering organization or team The SSCFM model illustrates that a software product integrates both first-party and third-party components, which are subsequently packaged for downstream use.
+These integration points also represent potential vulnerabilities where the software product is at risk of malicious code injection from either internal or external sources.
+Figure 1: The Software Supply Chain Factory Model [79].
 The study of software supply chains has been driven by engineering organizations’ greater reliance on external components.
-Many works have reported on this in various software contexts.
+Many works have reported on this reliance in various software contexts.
 For example, a 2024 Synopsys report stated that, in 1,067 projects across 17 industry sectors, $\sim$96integrated open-source components, and $\sim$77of total code is from open source [76].
-Even in government- and safety-critical systems, the use of third-party open-source components has moved from anathema to normal [23, 24].
+Even in government and safety-critical systems, the use of third-party open-source components has moved from anathema to normal [32, 22].
 Analyses of open-source software package registries have shown large increases year-over-year in the number of available components, the number of actively-used components, and the number of transitive dependencies [91, 99, 38, 54].
-When an external component contains a vulnerability, that vulnerability may be propagated to its dependents.
-The increased reliance on external components in software production has amplified the potential for exploitation of these vulnerabilities, and many attack scenarios have been proposed [29, 30, 31, 32, 33].
-For example, Lauinger et al.
+The increasing reliance on external components in software production amplifies the potential for the exploitation of vulnerabilities that may be contained in these components.
+As an example of this risk, Lauinger et al.
 [46] report that $\sim$37of the most popular Internet websites rely on at least one library with a publicly acknowledged vulnerability.
-Stock et al.
-showed that among 1,273 real-world client-side XSS vulnerabilities, 273 cases occurred exclusively within third-party code [35].
-Upstream in the supply chain, Zimmerman et al.’s analysis of npm indicated that up to 40% of all packages depended on code with at least one publicly known vulnerability [26].
-Given the prevalence of such inherited vulnerabilities, Zahan et al.
+Given the prevalence of inherited vulnerabilities, Zahan et al.
 [98] and similar studies have summarized attack vectors and potential weak links in software supply chains [25, 9, 71, 96, 52, 29, 56, 43].
@@
 Okafor et al.
-[45] characterize them as promoting three distinct dimensions: transparency (knowing all components in the chain), separation (isolating components), and validity (ensuring integrity of components).
+[57] state that these approaches promote three distinct dimensions: transparency (knowing all components in the chain), separation (isolating components), and validity (ensuring integrity of components).
 For example, transparency is increased by adoption of Software Bills of Materials (SBOMs) [51, 95].
-Separation is promoted by techniques such as sandboxing [48, 49, 50, 51, 52] and zero-trust architectures [53, 54].
-Validity is improved by reproducible builds [55, 56], software signatures [5], and attestations about artifact metadata such as authorship [57].
+Separation is aided by techniques such as sandboxing [74, 4, 45, 83, 82] and zero-trust architectures [73, 8].
+Validity is improved by reproducible builds [44, 86], software signatures [64], and attestations about artifact metadata (e.g., authorship) [81].
 Security standards, frameworks, and regulations have been established to complement and direct these security efforts.
-Notable examples include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [20], the Cloud Native Computing Foundation’s (CNCF) Software Supply Chain Best Practices [58], Microsoft’s Supply Chain Integrity Model (SCIM) [59], and the US National Institute of Standards and Technology’s (NIST) Secure Software Development Framework (NIST-SSDF) [60] and NIST 800-204D [61], which were developed in conjunction with US Executive Order 14028 (EO-14028) [62].
-At present, national regulations are inadequate to force compliance with secure supply chain practices [63].
+Notable examples include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [79], the Cloud Native Computing Foundation’s (CNCF) Software Supply Chain Best Practices [2], Microsoft’s Supply Chain Integrity Model (SCIM) [3], and the US National Institute of Standards and Technology’s (NIST) Secure Software Development Framework (NIST-SSDF) [18].
+Current national regulations do not force compliance with secure practices [48].
 2.2.2 Software Signing in Detail
 Software signing is a formally guaranteed method of establishing the authorship of a software component [28, 27, 66, 20].
-Signing ensures the provenance of components through public-key cryptography (Figure 2), improving validity (provenance) in supply chains [45].
-Many standards and government regulations recommend signing as a base Software Supply Chain security technique [20, 7, 67, 59, 9].
-For example, in The Linux Foundation’s SLSA framework, software signing is needed for Build Security Level 3 [20].
-Figure 2: Typical workflow for software signing and verifying signatures.
-The component author packages (A) and signs (B) their software.
-The signed package (C) and public key (D) are published.
-To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
-2.3 Empirical Work on SW Supply Chain Security
-The theoretical properties of techniques to improve software supply chain security are well known, whether via transparency, separation, or validity.
-However, adopting them in practice remains a challenge.
-Empirical/in vivo studies on software supply chain security have examined adoption practices and challenges to identify gaps in theoretical guarantees or usability.
-Some studies examine the general activities of component management, e.g., studies on dependency selection [68] and maintenance [69], and examinations of security challenges and incident handling procedures [70].
-Others have examined specific techniques for supply chain security.
-For example, with respect to the transparency property, researchers have studied the practices of SBOM adoption [71, 72], and reported gaps in tooling and interoperability as well as incomplete adoption and maintenance across the software development lifecycle.
-With respect to software signing, researchers have examined gaps in usability and adoption.
-Older implementations of signing were complex to use [73, 74], inspiring newer approaches with simplified tooling and support for key management [75, 76, 77].
-About 25% of publicly-reported attacks from 2010-2021 exploited the missing or erroneous use of signing [15].
-Gokkaya et al.
-showed that the top signing-related attack vectors were self-signed code, broken signature systems, and stolen certificates [42].
+It ensures the provenance and integrity of components in the supply chain through public-key cryptography.
+The process involves creating a hash of the software, encrypting it with the author’s private key, and attaching the resulting signature to the software.
+When users download the software, they decrypt the signature with the corresponding public key and compare the hash to ensure the software has not been altered.
+A typical workflow for a package signing is shown in Appendix A.
+Keyless signing[53, 35] eliminates the need for authors to manage long-term private keys.
+Instead, authors authenticate with an identity provider, and an ephemeral key associated with their verified identity is issued for each signing.
+Users then additionally verify that the public key used for the software verification is associated with the intended author, ensuring both software integrity and authenticity.
+Many standards and regulations recommend signing as a base Software Supply Chain security technique [79, 80, 70, 3, 17].
+For example, in The Linux Foundation’s SLSA framework, software signing is part of Build Security Level 3 [79].
+2.3 Empirical SW Supply Chain Security
+Although many theoretical techniques for software supply chain security are known, adopting them in practice remains a challenge.
+Empirical/in vivo studies on software supply chain security have examined adoption practices and challenges to identify gaps in theoretical guarantees or usability[90, 62, 29].
+Some studies examine the general activities of component management, e.g., dependency selection [60] and maintenance [11], and examinations of security challenges and incident handling procedures [88].
+Others have examined techniques for supply chain security.
+For example, for the transparency property  section 2.2.1, researchers have studied the practices of SBOM adoption [94, 41] and reported gaps in tooling, interoperability, adoption, and maintenance.
+For separation, studies have examined the performance of sandboxing [87] and its adoption in open-source ecosystems [7].
+For validity, reproducible builds have been analyzed in PyPi [86].
+Concerning software signing, researchers have examined challenges in its usability and adoption that create opportunities for attacks [90, 62].
+In 2010–2021, roughly 25% of publicly reported attacks exploited the missing or erroneous use of signing  [36].
+Top signing-related attack vectors include self-signed code, broken signature systems, and stolen certificates  [29].
+Although software signing is valued, it is considered costly to set up, as reported in Ladisa et al.’s survey  [43, 89].
 Several reports show that signing adoption rates for open-source components are low.
-Zahan et al.
+For instance, Zahan et al.
 found that $\leq$0.5of packages in the npm and PyPI registries had signed releases  [97].
@@
 corroborated this proportion, further noting that signing adoption is correlated with registry policies and that the correctness or quality of signatures depends on usable tooling  [64].
-Ladisa et al.’s survey shed light on these measurements, finding that software signing is valued but is considered moderately costly to set up [44, 78].
 3 Knowledge Gaps & Research Questions
@@
   • RQ1:
-    How and why is software signing implemented by software teams?
+    Where and how is software signing implemented by software teams?
   • RQ2:
@@
   • RQ4:
-    How does software supply chain failures and standards influence the adoption of software signing?
+    How do internal and external signing events influence the adoption of software signing?
 4 Methodology
@@
 We depict our methodology in Figure 2.
-In this section we discuss our research design rationale (section 4.1), protocol development (section 4.2), participant recruitment (section 4.3, data analysis (section 4.4), and study limitations (section 4.5).
+We describe our research design (section 4.1), protocol development (section 4.2), participant recruitment (section 4.3), data analysis (section 4.4), and limitations (section 4.5).
+Our Institutional Review Board (IRB) approved this study.
 Figure 2: Study methodology. Several sources (left) informed our protocol topics (center). We applied thematic and framework analyses (right).
-4.1 Approach — Research Design Rationale
-Our questions require insight into practitioner perspectives.
+4.1 Research Design and Rationale
+Our research questions require insight into practitioner perspectives.
 We considered two factors in selecting a methodology.
-First, since the prior work in this space is limited, there was relatively little basis on which to structure a closed-ended survey.
-Second, we expected recruitment to be a problem, since the target population has high security expertise and a broad organizational view, and thus are likely to be high-ranking (and busy) people.
-We therefore opted for a semi-structured interview [79], which would permit us to capture emergent results [80] (factor 1), and to leverage our professional relationships to improve response rates (factor 2).
+First, because the prior work in this space is limited, there was relatively little basis on which to structure a closed-ended survey.
+Second, our research questions warrant long-form descriptive-type answers that have to be extracted through open-ended questions, opinions, and detailed observations.
+We therefore opted for a semi-structured interview [63], which would permit us to capture emergent results [21].
 The resulting data is expected to be exploratory, and rooted in practitioners’ perspectives and experiences [63].
@@
 We therefore adopt an interpretivist perspective for our research methodology design and analysis.
-An interpretivist approach treats each situation as unique [82], focusing on in-depth variables and factors related to each context [83].
+This approach treats each situation as unique [10], focusing on in-depth variables and contextual factors [6].
+Our choice of the interpretivist framework was also motivated by our sampling technique as we discuss later (section 4.3).
 4.2 Protocol Design and Development
-To create the initial interview protocol, we studied three kinds of resources.
-First, to ensure a common language and a comprehensive set of topics, we studied software supply chain security frameworks such as SLSA, CNCF, SCIM, and NIST-SSDF (section 2.2).
+We studied three kinds of resources as we created the initial interview protocol.
+  1. 1.
+    Software Supply Chain Security Frameworks: To ensure a common language and a comprehensive set of topics, we studied software supply chain security frameworks such as SLSA, CNCF, SCIM, and NIST-SSDF (section 2.2).
 These frameworks use variants of the Software Supply Chain Factory Model (SSCFM, cf.
@@
 This choice allowed us to investigate signing practices at various stages of the SDLC while using a model familiar to our subjects.
-Second, we examined grey literature sources to gain insights into some practical uses of software signing.
-Our primary reference was the official Sigstore blog as of Jan.
-31, 2023.
-This resource has eight user reports about adopting software signing in various industry contexts, from which we derived questions about the contextual and organizational factors that subjects consider.
-Third, we reviewed adjacent literature [68, 69, 84] to understand effective questions related to software supply chain security, while avoiding duplication.
+  2. 2.
+    Grey Literature: To gain insights into practical uses of software signing, we examined grey literature such as industry blogs [30, 67], whitepapers [16], and case studies [23, 58, 93].
+These sources documented the adoption of software signing in various industry contexts, and, from these, we derived questions about the contextual and organizational factors that subjects consider.
+We collected these grey literature sources through Web searches of software signing tools mentioned or recommended by any of the security frameworks we studied.
+Additionally, we reviewed linked articles related to software signing in these standards.
+  3. 3.
+    Background & Literature Review: Our background and review of related works (section 2) provided a foundation for knowing what questions would be effective for answering our research questions while avoiding duplication.
+Broadly, our literature review process, which yielded all works cited in section 2) was based on a forward and backward snowball search.
+We began with works published in prominent cybersecurity and software engineering venues (USENIX, CCS, IEEE S&P; ICSE, FSE, ASE) and covered a 10-year period.
 After protocol development, we refined the protocol through practice interviews (2) and pilot interviews (2).
 For practice, the lead author interviewed two of the secondary authors.
-These practice sessions led to modifications in sections B and C of our interview protocol (see Appendix Appendix C).
+These interviews led to modifications in sections B and C of our interview protocol (see Appendix F).
 After recruiting subjects, we treated the first two subjects as pilot interviews.
 During these pilots, we gathered additional insights into how subjects perceived software supply chain attacks and reorganized the protocol, particularly focusing on questions related to the perceived sources of greatest risk (section B).
-Since few questions were changed or added, we included the data from the pilot interviews in our study where feasible.
-The final change occurred after the seventh interview.
-The author team discussed the results so far and identified one emergent topic: how practitioners’ organizations integrate feedback into their SSC security processes.
-One question was added for use in the remaining interviews.
-Our interview protocol is summarized in Table I, with a full version in Appendix Appendix B.
-TABLE I: Summary of interview protocol. See Appendix Appendix B for full protocol. Topic D was not analyzed in this study.
+Since few questions were changed or added, we included the data from the pilot interviews in our study where feasible  [^1] We included data from the pilot interviews in our analysis, as qualitative methods guidance supports this practice when the protocol demonstrates its ability to effectively answer research questions  [78, 37].
+This approach is particularly justifiable in contexts where recruitment is challenging, as was the case here and in similar work  [49].
+Quantitatively, we observed that: (1) the pilot interviews used nearly all ( 92%) of the final protocol, as only 3 out of 35 questions were changed afterward, and (2) the unique codes derived from the pilot interview transcripts were comparable in quantity and richness to those from other participants (see  Figure 5)..
+The final change was made after the seventh interview.
+The author team reviewed the results up to that point and identified an emergent topic: how practitioners’ organizations integrate feedback into their software supply chain (SSC) security processes.
+Consequently, one additional question was included in the remaining interviews, to address this emerging topic.
+Defining Internal & External Events to Answer RQ4: Unlike the other research questions, RQ4 focuses on investigating how internal and external factors influence organizational decisions to adopt software signing.
+Following the definitions of Schorlemmer et al.
+[64], we scope this part of the protocol to the impact of (1) Software Supply Chain failures, and (2) Software Supply Chain standards.
+Table 1 summarizes our interview protocol. The full protocol is given in Appendix E.
+Table 1: Summary of interview protocol. Topic D is not analyzed in this paper. See Appendix E for full protocol.
 Topic (# Questions) Sample Questions
@@
 4.3 Data Collection
-Our Institutional Review Board (IRB) approved this study.
-TABLE II: Subject Demographics.
+Table 2: Subject Demographics.
 For anonymity, we used generic job roles, not specific titles.
-“Senior management” refers to senior managers, directors, and executives.
-“Technical leader” — senior, lead, partner, and principal engineers.
-“Engineer” and “Manager” — junior staff.
-ID Role Experience Product Area
-S1 Research leader 5 years Digital technology
-S2 Senior mgmt. 15 years SSC security
-S3 Senior mgmt. 13 years SSC security
-S4 Technical leader 20 years Cloud, Dev tools
-S5 Engineer 2 years Cloud security
-S6 Technical leader 27 years Digital technology
-S7 Manager 6 years Cloud security
-S8 Technical leader 8 years Internet services
-S9 Engineer 2.5 years SSC security
-S10 Engineer 13 years Dev tools
-S11 Technical leader 16 years Social technology
-S12 Technical leader 4 years Cloud, OSS security
-S13 Senior mgmt. 16 years Telecommunications
-S14 Research leader 13 years Cloud security
-S15 Senior mgmt. 15 years Aerospace security
-S16 Technical leader 26 years Digital technology
-S17 Senior mgmt. 15 years SSC security
-S18 Manager 11 years Internet services
-Target Population: To address our research questions effectively, the target population was industry practitioners with expertise in software signing.
-An ideal subject would have in-depth knowledge and experience related to software supply chain practices within their organizations.
-Such experts would have high credibility, insider perspectives, and nuanced interpretations.
-Recruitment: We employed a non-probability-based Purposive and Snowball population sampling approach to recruit such subjects [82].
-Our initial subject sample pool was members from the Cloud Native Computing Foundation’s Kubecon 2023 conference organizing committee, many of whom also contribute to software supply chain security projects hosted by The Linux Foundation.
-This pool contained many security experts, although we acknowledge that the connection to specific signing projects may bias their responses to this genre of signing approach.
-Subsequently, we expanded our subject pool through recommendations from this initial group of subjects and other industry contacts.
-We offered each participant a \$100 gift card as an incentive.
+“Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff.
+We also highlight type of software subject’s immediate team produces.
+* refers to cases where subject’s team built other types of software in addition to the main stated.
+O refers to cases where subject belonged to multiple teams with other products to the one stated.
+ID Role Experience Software Type
+S1 Research leader 5 years Internal POC software
+S2 Senior mgmt. 15 years SAAS security tool*O
+S3 Senior mgmt. 13 years SAAS security tool*O
+S4 Technical leader 20 years Open-source tooling
+S5 Engineer 2 years Internal security tooling
+S6 Technical leader 27 years Internal security tooling*O
+S7 Manager 6 years Security tooling*O
+S8 Technical leader 8 years Internal security tooling*
+S9 Engineer 2.5 years SAAS security*
+S10 Engineer 13 years SAAS security*
+S11 Technical leader 16 years Firmware*
+S12 Technical leader 4 years Consultancy
+S13 Senior mgmt. 16 years Internal security tool
+S14 Research leader 13 years POC security Software*
+S15 Senior mgmt. 15 years Internal security tooling
+S16 Technical leader 26 years Security tooling
+S17 Senior mgmt. 15 years SAAS
+S18 Manager 11 years Security tooling
+Target Population: To effectively address our research questions, the target population was industry professionals with expertise in software signing.
+An ideal subject would have deep knowledge and experience related to software supply chain practices within their organizations.
+Such subjects are typically middle- or high-ranking engineering staff.
+Recruitment: Given the seniority and expected workload of the target population, we anticipated recruitment to be a challenge.
+To mitigate this, we employed a non-probability-based purposive and snowball sampling approach to recruit participants [10].
+This approach incurs the risk of bias, but has the reward of access.
+Given the understudied topic at hand, we felt that the reward outweighed the risk.
+Our initial sample pool of subjects was made up of members of the Kubecon 2023 conference organizing committee of the Cloud Native Computing Foundation.
+This group matches our target population: all of them either worked as a member of the software security teams of their organizations or worked in organizations with products and services in the domain of software security, and all were medium- to high-ranking members of their organizations.
+Many of these individuals also contribute to software supply chain security projects hosted by The Linux Foundation, making them well-suited to provide expert insights.
+We had special access to this group because one of the authors was a member of that committee.
+However, recognizing the potential biases of this initial sample, we expanded our recruitment through recommendations from initial participants and other industry contacts.
+This approach allowed us to diversify our subject pool and gather a broader range of insights, while still focusing on individuals with substantial expertise in the domain.
 In total, we contacted 30 candidates, of whom 18 agreed to participate (60% response rate).
 These subjects came from 13 different organizations.
-Their demographic information is summarized in Table II.
-Subjects were affiliated with companies with security products or were part of their company’s security team.
-For subject anonymity, we do not distinguish those from CNCF or from snowballing.
+The subjects were affiliated with companies with security products or were part of their company’s security team.
+Each participant was offered a \$100 gift card as an incentive, in recognition of their high rank.
+Subject Demographics: We summarize the demographic and organization information of the subjects in Table 2 and  Table 6.
+We note that six participants were from the initial KubeCon committee invitations; industry contacts of the authors yielded seven participants; and snowball sampling (recommendations from participants) yielded five participants.
+In addition to the highlighted, we also note that 13 subjects (from 11 organizations – 4 Small, 2 Medium, and 6 Large organizations) confirmed that they either lead, or belong to, teams involved in software supply chain, infrastructure and tooling, or security control implementation.
+These subjects indicated that they are responsible for initiating or implementing their organization’s security controls and have responsibilities related to compliance.
+Although not all subjects directly owned compliance, their involvement in security strategy and operational decision making provided them with the appropriate context to assess the utility, influence behind adoption, and challenges affecting the implementation of Software signing.
+Subjects from small organizations typically focused on SAAS products related to software supply chain, container, and cloud security; or were consultants.
+In medium-scale organizations, subjects had more specialized roles, e.g., developing internal security-enabling software for infrastructure.
+In large organizations, subjects had similar goals, but often belonged to/oversaw multiple teams.
+For confidentiality, we refer to organizations by the letters A–L and participants using the notation SX(y), where $X$ denotes a unique subject number and $y$ either distinguishes participants from the same organization or represents the subject’s product type.
+This notation is also used to provide context to subject’s responses.
+For example, S10 (SAAS, Dev tools) refers to Subject S10’s direct team’s software product, which is a SaaS-type product for an organization whose product area is developer tools.
+Addressing Potential Bias: While the connection to the CNCF and specific signing projects could bias responses towards certain approaches, this potential bias is balanced by the deep involvement of participants in the field of software supply chain security.
+Their contributions are crucial for understanding the current state and future directions of software signing practices.
+Moreover, the insights provided by these experts, who are at the forefront of the industry, offer a unique and valuable perspective that is essential for the study’s objectives.
+Generalizability of Results: We acknowledge that the non-random sampling method and the concentration of participants from a specific initiative may limit the generalizability of our findings to the broader population of industry practitioners.
+However, the purpose of this study was not to achieve broad generalizability but rather to gain in-depth insights from a highly specialized group of experts—hence the use of an interpretivist analysis framework.
+The findings are particularly relevant to organizations and stakeholders involved in similar software supply chain security initiatives and may inform best practices and future research in this specific context.
+While broader generalization may be limited, the detailed, expert-driven analysis provided by this study offers significant value in advancing understanding within this niche but critical area.
 Interviews: Interviews were conducted by the lead author via Zoom. The median interview duration was 50 minutes.
 4.4 Data Analysis
-Interview recordings were transcribed by www.rev.com (human transcription). We anonymized identifying information. Our analysis then proceeded in several stages.
-(Stage 1) We began with data familiarization [85]. Three analysts (A1–A3) undertook an inductive reading, memoing, and coding process for all transcripts.
-(Stage 2) Subsequently, following O’Connor & Joffe’s recommendation [86] to initially independently code 10-25% of the data, we randomly selected approximately 33% (6 transcripts) for multiple memoing.
-Two analysts (A1, A2) conducted this multiple memoing of data in three rounds, collaborating to discuss emerging code categorizations after each round.
-This process yielded 506 coded memos, which were reviewed by Analysts A1 and A3 over four one-hour meetings to create a codebook comprising 36 code categories, excluding demographics.
-(Stage 3) To ensure the reliability of our coding, analysts A1 and A3 randomly selected and coded a transcript.
-We used the Percentage agreement metric as recommended by Feng [87] for easy coding tasks with nominal data.
-Since our codebook was collectively defined by both analysts A1 and A3, we assumed this coding task should not be too difficult — we achieved an 89% agreement rate, confirming this assumption.
-Any discrepancies in coding were resolved through discussion and code redefinition.
-(Stage 4) Next, a second set of 6 transcripts was selected and coded by Analyst A1 using an initial codebook, adding 6 minor categories, each with less than 6 codes.
-This updated codebook was reviewed by Analyst A3.
-(Stage 5) The same process was applied to the final 6 transcripts, yielding 4 more minor categories with fewer than 4 codes each.
-Out of the 10 new categories, only 2 provided new insights into our data, as determined through multiple review rounds by Analysts A1 and A3.
-(Thematic and Framework Analyses) From this process we induced themes following Braune & Clark’s method [88].
+Interview recordings were transcribed by www.rev.com (human transcription).
+We anonymized identifying information.
+Our analysis then proceeded via multi-stage coding  [13] and a subsequent thematic  [12] and framework analysis  [72].
+Three analysts (A1–A3) participated in the analysis stage of the transcripts.
+Analyst A1, the lead author, was responsible for all stages of data collection (including protocol design and interview conduct) and contributed to all phases of the analysis.
+Initial analysis was conducted by A1 and A2.
+When A2 became unavailable, we recruited A3.
+To ensure consistency, A3 was trained by memoing and coding transcripts that had already been analyzed.
+Comparing A3’s memos to A1 and A2, we observed strong alignment, ensuring reliability in the analysis process.
+A1-A3 are graduate students with research experience in cybersecurity, including software signing.
+The supervisors are researchers in software engineering and cybersecurity.
+Multi-stage coding:
+Stage 1. We began with data familiarization [77].
+Stage 2.
+Next, we needed to develop and refine our codebook.
+To provide reliability while conserving resources, we used the technique described by Campbell et al.
+[13] and O’Connor & Joffe  [59].
+First, we chose a proportion of random transcripts to multiply-code (i.e., multiple analysts coding).
+[^2] Although there is no common agreement on a suitable proportion of the data set to select for this, O’Connor and Joffe recommend 10–25% as typical [59].
+We randomly selected approximately 33% (6 transcripts) to develop our codebook.
+This decision to slightly exceed O’Connor and Joffe’s recommended percentage was to familiarize analysts A2 and A3 with the nature of our collected data, proposed memo, and coding formats.
+Then, two analysts (A1, A2) conducted multiple rounds of memoing on these transcripts, discussing emerging code categorizations after each round.
+This process resulted in 506 coded memos, which were reviewed by Analysts A1 and A3 in a series of meetings.
+Ultimately we obtained a codebook with 36 code categories.
+This codebook was further refined and extended as discussed in Stages 3-5.
+Stage 3.
+To assess the reliability of our coding, we combine the techniques of Campbell et al.
+[13] and Maxam&Davis [50].
+Analysts A1 and A3 randomly selected and coded a transcript independently [13, 55], and we measured their degree of agreement.
+Since our codebook was collectively defined by analysts A1 and A3, we supposed this coding task would not be difficult, and thus we used the percentage agreement metric as recommended by Feng [26].
+In this independent coding step, we observed an 89% agreement rate.
+The analysts were able to resolve disagreements through discussion and code refinement.
+According to Campbell et al.
+[13], this high level of agreement suggests that a coding scheme is suitable for a single coding of the remaining uncoded transcripts.
+Stage 4.
+Next, a second set of 6 transcripts was selected and coded by Analyst A1 using the resultant codebook from Stage 3.
+Following Campbell et al.’s recommendation [13], we created new code categories as they emerged; this phase added 6 new code categories, each with less than 6 associated coded memos.
+Analyst A3 reviewed this updated codebook.
+Stage 5.
+The same process was applied to the final 6 transcripts, yielding 4 more code categories with fewer than 4 coded memos each.
+Of the 10 new categories recorded from stage 4 and 5, only 3 provided new insights into our data, as determined through multiple review and discussion rounds by Analysts A1 and A3.
+Thus A1 and A3 agreed that these new codes be classified as minor code categories.
+Thematic and Framework Analyses: From this process, we induced themes following Braune & Clark’s method [12].
 These themes helped us answer RQs 2-4 directly.
-To answer RQ1, we also applied a framework analysis [89], mapping the themes to the software supply chain factory model as a reference framework to analyze participants’ responses and further categorize the themes.
-Through our coding phase’s thematic analysis, we identified eight themes.
-Four of these themes were focused on the practical implementation of Signing, while the remaining four addressed various aspects of software signing: the challenges encountered, its perceived importance, the impact of security failures on its adoption, and the influence of security standards and regulations on its adoption.
+To answer RQ1 we also applied a framework analysis [72], mapping the themes to the software supply chain factory model as a reference framework to analyze participants’ responses and further categorize the themes.
+This framework was selected because it serves as the general reference for SSC security.
+Through our coding phase’s thematic analysis, we identified eight themes  Figure 6.
+Four of these themes were focused on the practical implementation of software signing, while the remaining four addressed various aspects of software signing: the challenges encountered, its perceived importance, the impact of security failures on its adoption, and the influence of security standards and regulations on its adoption.
 To determine whether additional aspects of the phenomenon might be expected to emerge if we continued to sample, we measured saturation in the 18 interview transcripts.
-Following Guest et al.’s [90] recommendation, we tracked the cumulative appearance of new codes in each interview.
-Each interview had a median of 33 unique codes, also depicted in  Figure 5.
-Saturation was achieved after interview 11 (details in Appendix Appendix D).
-4.5 Study Limitations
-Positionality Statement: We acknowledge the role that our backgrounds may have played in this study [91].
+Following Guest et al.’s [33] recommendation, we measured saturation by tracking the cumulative appearance of new codes in each interview.
+Each interview was rich, with a median of 33 unique codes (Figure 5).
+As shown in Figure 5, saturation was achieved after interview 11.
+4.5 Limitation and Threats to Validity
+We discuss construct, internal, and external threats to validity [92].
+Following the guidance of Verdecchia et al.
+[84], we focus on substantive threats that might influence our findings.
+We also describe the role that our own perspectives and experiences may have played in this research (positionality) [24].
+Construct Threats are potential limitations of how we operationalized concepts.
+To mitigate these threats, we grounded our interview protocol in the academic and grey literature on software signing.
+Following best practices, we also piloted the interview protocol internally and externally to assess its utility in capturing the desired data [15].
+Internal Threats are those that affect cause-effect relationships.
+Qualitative data implies subjective analysis.
+To improve the reliability of our results, we followed an iterative process to develop the codebook and used multiple raters with a high measured level of agreement.
+We also acknowledge the possibility of a response bias, as some participants were recruited from the professional network of one of the authors.
+To address this, the involved author was not present during the interviews and played no role in the data analysis process.
+External Threats may impact generalizability.
+This work is qualitative, which inherently carries limitations.
+Given the cost of an interview study, our work has a smaller sample size (N=18) than methods such as surveys can achieve.
+Our sample size is consistent with that of similar studies [94, 41].
+Our subject recruitment approach may also have introduced biases, e.g., towards CNCF/The Linux Foundation-oriented frameworks and tools.
+As a trade-off, using CNCF members as a starting point improved our access to high-ranking engineering staff.
+To reduce bias, we also recruited through snowball sampling and other professional networks.
+We assessed this risk with saturation — by the 11th interview our data saturated, supporting the adequacy of our sample size (Figure 5).
+We also segmented our interview protocol (Table 1) and intentionally excluded the tool-specific analysis from this study.
+Instead, we concentrated on parts of the protocol related to the general practice of software signing.
+Our study focuses on an expert population to capture nuanced insights into software signing strategies and decisions.
+While this approach excludes the perspectives of the broader developer community, it aligns with our objective to understand signing practices at a strategic level.
+Less experienced developers may interpret or prioritize signing differently, reflecting their vantage point within the software development lifecycle.
+By targeting experts, we ensured that the findings are informed by those directly responsible for implementing and guiding signing strategies.
+Positionality Statement: We acknowledge that our backgrounds may have influenced this study [24].
 The author team are cybersecurity researchers with expertise in software signing, so our expertise both enabled and influenced the protocol development and analysis.
-Beyond this, one of the authors is a contributor to a popular signing tool.
-To reduce bias in our study, that author had no involvement in the development of the research protocol or analysis.
-However, that author reviewed the results and framing of the study, to increase its validity for the software signing community.
-Other Threats to Validity: This work is qualitative, and bears the limitations of any qualitative study.
-First, the cost of data collection reduces the sample size, threatening generalizability.
-Our subject recruitment may also have introduced biases, e.g., towards CNCF/The Linux Foundation-oriented frameworks and tools.
-As mitigations, we note that our codebook saturated at the 11th interview, with few/no new codes observed in subjects 12-18 (Appendix D); and our sample size is typical for similar studies [71, 72].
-Second, qualitative data implies subjective analysis.
-To improve the reliability of our results, we followed an iterative process to develop the codebook and used multiple raters with a high measure of agreement.
-Specific to the topic of software signing, the frameworks for software supply chain security and the available automation for implementing signing are evolving.
-Our results should be understood as a snapshot of current practices in industry.
-They will grow out of date as signing standards, regulations, and tooling continues to evolve.
+In addition, one of the authors is a contributor to a popular signing tool.
+To reduce bias resulting from specific aspects of that tool, that author had no involvement in the development of the research protocol or analysis.
+However, that author informed the framing of the study and reviewed the results.
 5 Results
 5.1 RQ1: Signing Implementations in Practice
-Finding 1.
-We refine the software supply chain factory model from Figure 1, as shown in Figure 4 and Table III.
-Our refined model shows candidate points for the establishment of provenance.
-We analyze our subjects’ responses using this refined model, identifying points on this model where subjects’ teams or organizations employ software signing.
-Figure 4: Our refined software supply chain factory model (compare to Figure 1), showing different points where provenance and integrity of software could be established (PI, PE, PB, PD, PS, PA) and Verified (VI, VE, VB, VD, VS, VA) using software signatures.
-Abbreviations are defined in Table III.
-TABLE III: Different Signing use Points as noted in  Figure 4 and subjects who employ software signing at these points.
-We mark subjects who consider them optional for reasons of demand by customers, nature of products, etc.
-Point Description Subjects
+Part of our interview protocol (Table 1) ascertained how subjects use software signing in their security strategies at each stage of the software supply chain factory model (Figure 1).
+From these responses, we identified when Software signing was required by the subject’s team (or organization) and mapped these instances back to the stages of the software supply chain factory model.
+The summary of our framework analysis is presented in Table 3.
+Table 3: Different Signing Use Points, categorized by the stages of the software supply chain factory model highlighted in Figure 1 and subjects who employ software signing at these points.
+We mark subjects who consider them optional for reasons of demand by customers, nature of products, team requirements, etc.
+The points highlighted (PI, VI, PE, VE, PS, VS, PB, VB, PA, VA, PD, VD) are described in  Figure 3.
+Points & description of Software signing use Subjects
 SOURCE CODE
-PI Internal code contributors sign their commits S2-S11, S14-S17 (S4, S8, S14 – Optional)
-PE External code contributors sign their commits S7, S10 (S11, S14 — Optional)
-VI Verification of signatures from Internal contributors S7, S14, S17
-VE Verification of signatures from external contributors
-PSC Signing after code reviews/audits S2-S11, S14-S17 (S4, S8, S14 — Optional)
+Internal code contributors sign commits-PI S2-S11, S14-S17 (S4, S8, S14–Optional)
+External code contributors sign commits-PE S7, S10 (S11, S14–Optional)
+Verify signatures from Internal contributors-VI S7, S14, S17
+Verify signatures from external contributors-VE
+Signing after code reviews/audits-PS S2-S11, S14-S17 (S4, S8, S14–Optional)
 BUILD
-VSC Verification of source code signatures before build S2, S3, S7, S14
-PB Signing of build output S1-S5, S7, S8, S11-S15, S17 (S8 —Optional)
+Verify source code signatures before build-VS S2, S3, S7, S14
+Signing of build output-PB S1-S5, S7, S8, S11-S17 (S8-Optional)
 DEPLOYMENT/PACKAGE/ARTIFACT
-VB Verification of build signature before Deployment or final artifact S7, S14
-PA Signing of Final Software product All S1–S18
+Verify build signature before deployment-VB S7, S14
+Signing of Final Software product-PA All S1–S18
 THIRD-PARTY DEPENDENCY
-PD Signing after verification and certification of dependencies S6, S7, S9, S16, S17, S18
-VD Verification of Signatures from external dependencies S9, S11, S17 (S9, S11, S17 — Optional)
-VA Verification by final customers Difficult to determine
-Using our initial analysis plus the recommendations and philosophy of several frameworks, which recommend a continuous verification of integrity and trust in code alone (not humans) [20, 92], we present a refined form of the software supply chain factory model as shown in Figure 4.
-Our model comprehensively identifies points in the software delivery process where the security property of provenance is expected to be established or verified.
-We compare this with subjects’ reported usage of other security methods for these factory model components, and second with framework (and standards) prescriptions.
+Signing after verification and certification of dependencies-PD S6, S7, S9, S16, S17, S18
+Verify signatures from external dependencies-VD S9, S11, S17 (S9, S11, S17 — Optional)
+Verification by final customers-VA Difficult to determine
 5.1.1 Source Code
-First, we ascertained whether our subjects used any form of external code contributors in their software creation process.
+First, we ascertained whether our subjects used any form of external code contributors in their software authorship process.
 Out of 10 subjects who reported a form of collaboration with open-source or other external contributors, only 5 subjects used external code contributions in the production code of projects they owned (both commercial and OSS).
-We then proceeded to determine if any form of signing is used to establish provenance for points PE, PI, VE, VI, and PSC.
-External contributors sign commits (PE): Signing is mostly required by teams who use external code contributors.
-Out of 5 subjects whose teams used external code contributions in their production code, 4 of these subjects require that commits be signed.
-However, two of those do not require signing for open-source projects or noncritical projects.
-Only 2 of our subjects unconditionally require signing from their external contributors.
-As S2 put it, “I got something signed from somebody I don’t know…So we do the code review, and then we have a sign-off by somebody on our team that signs off from that code”.
-Internal contributors sign commits (PI): Signing is almost always required by teams for their internal code contributors.
-14 of 18 subjects report that their team required some sort of commit signing from internal contributors.
-Subjects who list this as an optional requirement (3) cite similar reasons as in PE above.
-Verification of signatures from contributors (VE&VI): While signing is relatively required, it is not verified as a check for security and provenance of code contributions from contributors.
-Techniques like code reviews (and security audits), security/vulnerability scans, tests, etc, are the most used security methods for assessing the security of code contributions.
-S10 opined to this, “I think everybody on the team does sign their commits, but we don’t really verify it…If somebody stopped signing their commits, I don’t think much would really happen”
-Code is signed after reviews/audit (PSC): Signing here is common similar to PI.
-Subjects who report that signing is required in PI, also report that all activities related to code reviews, or audits are (cryptographically) signed off on.
+We highlight the points and manner in which subjects reported the use of software signing to create their source code.
+External contributors sign commits: Most teams who had input from external code contributors required Software signing.
+Out of 5 subjects (4 organizations) whose teams used external code contributions in their production code, 4 of these subjects (3 organizations) require that commits be signed.
+However, two of those do not require signing if the contribution is toward their open-source projects or noncritical projects.
+Only 2 of our subjects (2 organizations) unconditionally require signing from their external contributors.
+As S2 (SAAS, SSC Security tool) put it, “I got something signed from somebody I don’t know…So we do the code review, and then we have a sign-off by somebody on our team that signs off from that code”.
+Internal contributors sign commits: Signing is typically required by teams for their internal code contributors.
+14 of 18 subjects report that their team required commit signing from internal contributors.
+Three subjects (S4, S8, S14) described this as an optional requirement and largely agreed with S2’s reasoning that signing alone does not establish trust in identity.
+Notably, these subjects’ teams often rely on contributions from both internal and external contributors.
+Verification of signatures from contributors: Although code contributions are often required to be signed, the signatures are not always verified to ensure the security and provenance of the contributions.
+According to our subjects, the most commonly used security methods for assessing the security of code contributions include techniques like code reviews, security audits, vulnerability scans, and tests.
+For example, S10 (SAAS, Dev tools) states, “I think everybody on the team does sign their commits, but we don’t really verify it…If somebody stopped signing their commits, I don’t think much would really happen.”
+Code is signed after reviews/audit: The outputs of code reviews and audits are generally required to be cryptographically signed.
+Subjects whose teams(or organizations) mandate signing for code contributions by internal contributors also report that all activities related to code reviews and audits must be signed as well (14 subjects).
 5.1.2 Build
-Next, we extract from our subjects’ responses how signing is used in the build process of software products.
-For most of the subjects, the build and build signing (for subjects who do so) are automated.
-Our model expects that the source code to be built should be verified (VSC), and then when the build is successful we expect another round of provenance certification (PB).
-Verification of source code signatures before build (VB): Signatures from source codes are rarely verified before the commencement of the build process.
-Two subjects mentioned some sort of rigorous verification at this point.
-Alternate means of security at this stage are similar to the security techniques reported for the build process.
-Signing of build output (PB): 13 subjects reported the use of signing at this stage of the software delivery process.
-10 subjects reported signing is used alongside attestations, SBOMS, and other forms of metadata.
+Next, we examine our subjects’ responses regarding how signing is used in the build process of software products.
+For most subjects, the build process and the signing of build outputs (for subjects who do so) are automated.
+We expect that automation increases the use of software signing.
+Verification of source code signatures before build: Signatures from source code are rarely verified before the build process begins.
+Four subjects mentioned some form of rigorous verification at this stage.
+For those who do not use or verify signatures, alternative security techniques are regularly employed, such as automating build pipelines with workflow definitions, using build policies, implementing reproducible builds, and utilizing code scanning tools to ensure the security of inputs to the build process.
+Signing of build output: 14 subjects (11 organizations) reported signing at this stage of the software delivery process.
+10 subjects (6 organizations) reported signing alongside attestations, SBOMS, and other metadata.
 5.1.3 Package/Artifact
 Next, we analyze the use of signing at the final stage of deploying software products.
-Verification of build signature before deployment of final artifact (VB): Similar to the other verification steps, signatures from the build phase are rarely verified before the deployment.
-Only two subjects mentioned verification at this point.
-For example, S7 states, “We sign our code and then we also verify that it was built in our specific build system”.
-Other subjects mentioned that the security of the entire build process comprises techniques like an automated build pipeline with Workflow definitions, use of build policies, and Reproducible builds.
-These are complemented by code scanning tools as in the case of the source code stage.
-Signing of Final Software product (PA): All 18 subjects said that signing is done at the deployment stage to establish provenance for the organization on each of their products.
+Verification of build signature before deployment of final artifact: Similar to the other verification steps, signatures from the build phase are rarely verified before the deployment.
+Only two subjects (1 organization) mentioned verification at this point.
+For example, S7 (Security tools, Cloud security) states, “We sign our code and then we also verify that it was built in our specific build system.”
+Signing of Final Software product: All 18 subjects said that signing is done at the deployment stage to establish provenance for the organization on each of their final products.
 5.1.4 Third-Party Dependencies
-For third-party dependencies in the Software Supply Chain, we evaluated how software signing is used as a factor influencing dependency selection, and how software signing is used to verify the authenticity of dependencies in practice.
-Presence of a Signature as a factor for Dependency Selection: For most of our subjects, signatures on a third-party dependency do not affect their choice in selecting such a dependency.
-Only 3 of our subjects like to see this, treating signatures as a signal that the owner is security conscious.
-Other security factors that influence the selection of third-party dependencies as mentioned are: the use of the OpenSSF’s scorecard tool, the dependency’s known vulnerabilities, level of activity, licensing, and age.
-Verification of Signatures from external dependencies (VD): Signatures on third-party dependencies are rarely verified in practice.
-Only three subjects reported some form of verifying signatures from third-party dependencies.
-As S11 said, “If it is open source, then probably no…But if it is commercial off the shelf, we want to make sure those are signed.” However, most of our subjects reported several methods for assessing the authenticity of a third-party dependency.
+For third-party dependencies in the software supply chain, we assessed how software signing influences dependency selection as a security factor and how signatures are used to verify the authenticity of dependencies in practice.
+Presence of a Signature as a Factor for Dependency Selection: When they select third-party dependencies, most subjects do not consider whether it is signed.
+Only three subjects view signatures as an indicator that the package maintainer prioritizes security.
+Other security factors influencing the choice of third-party dependencies include the use of OpenSSF’s scorecard tool, known vulnerabilities, level of activity, licensing, and the age of the dependency.
+Verification of Signatures from external dependencies: Signatures on third-party dependencies are rarely verified in practice.
+Only three subjects (3 organizations) reported some form of verifying signatures from third-party dependencies.
+Signatures are mainly verified for commercial off-the-shelf software.
+As S11 (Firmware & testing software, Social technology) says, “If it is open source, then probably no…But if it is commercial off the shelf, we want to make sure those are signed.” However, most of our subjects reported several methods for assessing the authenticity and integrity of a third-party dependency.
 These methods include code and vulnerability scanning (e.g., Coverity), source rebuilding, internal mirroring of dependencies, and reliance on ecosystem internal checks (e.g., Go package manager’s checks).
-Other subjects perform internal reviews of dependencies, while some have a centralized dependency selection and management process, where dependencies are hosted and maintained by the organization.
-A few subjects noted that their organization discourages or forbids third-party dependencies.
-Signing after verification and certification of dependencies(PD): A few subjects reported that third-party dependencies are thoroughly reviewed using one of the aforementioned techniques before being signed off for use by a team member.
+Other subjects perform internal reviews of dependencies.
+Some have a centralized dependency selection and management process, where the organization hosts and maintains dependencies (5 subjects from 4 organizations).
+A few subjects noted that their organization discourages or forbids third-party dependencies (S13 & S8).
+In addition to relying on alternative methods to assess dependency security, most of our subjects echoed the sentiment that OSS dependencies are rarely signed and that the identity of individuals signing open-source dependencies does not inherently establish trust.
+Signing after verification and certification of dependencies: A few subjects reported that third-party dependencies are thoroughly reviewed using one of the aforementioned techniques before being signed off for use by a team member.
 Six of our subjects indicated that this process is followed in their software production workflow.
-For example, S17 described their team’s process like this: “We have a signed report that’s signed by us, and we trust us…that says, ‘Hey, I scanned this third-party dependency, did all these things.
+For example, S17 (SAAS, SSC Security tool) described their team’s process as this: “We have a signed report that’s signed by us, and we trust us…that says, ‘Hey, I scanned this third-party dependency, did all these things.
 I don’t have a signature for that, but I ran all these extra checks and then signed those checks.
-So now we have a record”’.
-In contrast to that attestation approach, S18’s team ingests third-party dependencies and signs the internal versions directly.
+So now we have a record.”’ In contrast to that attestation approach, S18 (Security tooling, Internet service) says their team ingests third-party dependencies and signs the internal versions directly.
+5.1.5 Refined Software Supply Chain Factory Model
+As a result of our framework analysis, we identified various points where software signing is utilized.
+We compared these to the descriptions in several Software Supply Chain security frameworks, especially the SSCFM (section 4.2Figure 1).
+We observed that some of these frameworks and standards are ambiguous, particularly in specifying where within the software supply chain factory model practitioners are required to establish provenance, whether by creating or verifying a signature.
+We therefore present a refined version of Figure 1 in Figure 3.
+This updated model clearly identifies the points where provenance should (could) be established, providing guidance for organizations and teams newly adopting software signing in their software creation process.
+For example, our refined model is compared with the SLSA framework  [79].
+Despite SLSA’s rich description of security recommendations and its focus on build levels, there is surface-level confusion about whether provenance requirements apply solely to the build stage or also to earlier stages such as collaboration.
+Figure 3: Our refined software supply chain factory model (compare to Figure 1), highlighting different points where provenance and integrity of software could be established (PI, PE, PB, PD, PS, PA) and Verified (VI, VE, VB, VD, VS, VA) using software signatures.
+These points were identified from participants’ responses.
+Note that while signing at PB and PA stages was commonly practiced (in green), verification at VD, VB and VE stages was minimally performed (in red).
+Abbreviations are defined in Table 3.
 5.2 RQ2: Challenges Affecting the Implementation of Software Signing in Practice
-Finding 2.
-Subjects identified several common themes as challenges faced in practice as obstacles to effectively implementing software signing.
-We categorize these broadly; 1.
-Technical Challenges: These include – key Management and Infrastructure issues, compatibility issues with existing systems and different architectures, and lack of signature verification 2.
-Organizational and human Challenges: These include – expertise in setting up Signing implementation, operationalization of the signing procedure, developer attitude, and Resources to set up the signing process
 Our next research objective was to understand the various challenges practitioners face while implementing software signing in practice.
 We grouped subjects’ responses under three main themes for challenges faced by practitioners in practice; Technical, Organizational, and Human factors.
-Table IV gives a summary.
+Table 4 gives a summary.
 Due to space limitations, we focus on only a subset of the individual challenges.
-TABLE IV: Challenges to software signing Implementation in practice. We have categorized related challenges highlighted by subjects. Two uncommon Technical challenges are described in Appendix E.
-Observed Challenges #Subjects(#Orgs) Proposed Solutions
+Table 4: Challenges to software signing implementation in practice.
+We have categorized related challenges highlighted by subjects.
+Additional technical challenges, ’Managing Signed Artifacts’ and ’Identity Management/Authentication,’ are detailed in Appendix D, as they are not exclusive to signing tools.
+Observed Challenges #Subjects #Orgs Subjects’ Proposed Solutions
 Technical
-Key management 10 (9) Use of Keyless Signing (e.g., Sigstore)
-Compatibility issues 6 (6)
-Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
-Ease of use/Usability 4 (4) Usable signing tools (e.g., Sigstore), Documentation
-No unifying Standard 2 (2)
-Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
+Key Management 10 9 Use of Keyless Signing (e.g., Sigstore)
+Compatibility Issues 6 6
+Lack of Verification of Signatures 6 5 Signed Metadata, Component Data Management
+Ease of Use/Usability 4 4 Usable Signing Tools (e.g., Sigstore), Documentation
+No Unifying Standard 2 2
 Organizational
-Operationalization of the Signing process 4 (4) Automating Signing
-Resource to Set up Signing 3 (3)
-Creating Effective Signing Policy 2 (2) Regular Process Feedback mechanisms
-No Management incentive to Sign 2 (2)
-Bureaucracy 1 (1)
+Operationalization of the Signing Process 4 4 Automating Signing
+Resources to Set up Signing 3 3
+Creating Effective Signing Policy 2 2 Regular Process Feedback Mechanisms
+No Management Incentive to Sign 2 2
+Bureaucracy 1 1
 Human
-Expertise in setting setup and use of signing 5 (4)
-Developer Attitude to Signing 3 (3) Automating Signing
-Lack of Demand from Customers 1 (1)
+Expertise in setting setup and use of signing 5 4
+Developer Attitude to Signing 3 3 Automating Signing
+Lack of Demand from Customers 1 1
 5.2.1 Technical Challenges
@@
 Some subjects discussed recent signing technologies that implement keyless signing (e.g., Sigstore) as a possible solution.
-S18 said: “Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have these big servers to store all these signatures”.
-S17 also noted, “[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada.
-So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better.”.
-Compatibility issues: Compatibility issues arise during the integration of software signing systems and tools with other tools within the software engineering process, including other security tools and the artifacts to be signed.
-S2, S7, S17, and S18 highlight issues surrounding the compatibility of newer signing systems with existing software systems in the software engineering process.
-As S18 put it, “There are teams who’ve been doing this for quite a while…Selling that is a big challenge, right?
+S18 (Security tooling, Internet service) said: “Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have these big servers to store all these signatures”.
+S17 (SAAS, SSC Security tool) also noted, “[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada.
+So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better”.
+Compatibility Issues: Compatibility issues arise during the integration of software signing systems and tools with other tools within the software engineering process, including other security tools and the artifacts to be signed, they also include scenarios where challenges with signing were linked to other security toolings,.
+S2, S7, S17, and S18 highlight issues in the compatibility of newer signing systems with existing software systems in the software engineering process.
+As S18 (Security tooling, Internet service) put it, “There are teams who’ve been doing this for quite a while…Selling that is a big challenge, right?
 Because we need to prove that it is much more secure than what has been there in place.”
-S8, whose organization’s software product is built for different platforms and hardware architectures, noted compatibility issues across various architectures: “Every platform evolved separately…macOS…Windows…The requirements are changing and evolving independently”.
-Compatibility with other security tools can pose challenges too, as noted by S5: “The vulnerability scan report [is for]…code…CVE published…[but] signatures, I don’t think it will pick those things up.”
+S8 (Internal security tool & cloud APIs, Internet services), whose organization’s software product is built for different platforms and hardware architectures, noted compatibility issues across various architectures: “Every platform evolved separately…macOS…Windows…The requirements are changing and evolving independently”.
+Compatibility issues can also occur in situations where current security test tools are not designed to identify issues with Software signing, as noted by S5 (Internal security tool, Cloud security): “The vulnerability scan report [is for]…code…CVE published…[but] signatures, I don’t think it will pick those things up.”
 Ease of Use/Usability: Usability of the signing tool or implementation was a common challenge hampering effective use in practice.
-This was often expressed as developer experience, e.g., S17 said: “For a developer, make sure that they can sign the commits easily”.
-This point was also raised at a team level by S13: “I think that awareness of the value of signing and sophistication around how signing is performed has increased significantly, especially in recent years.
+This was often expressed as developer experience, e.g., S17 (SAAS, SSC Security tool) said: “For a developer, make sure that they can sign the commits easily”.
+This point was also raised at a team level by S13 (Internal security tool, Telecommunication): “I think that awareness of the value of signing and sophistication around how signing is performed has increased significantly, especially in recent years.
 But it has also resulted in a lot of smaller teams deciding that this is not a manageable thing.”
-With respect to documentation, S11 said: “I think there are multiple challenges there.
+Concerning documentation, S11 (Firmware & testing software, Social technology) said: “I think there are multiple challenges there.
 When you build some signing infrastructure, you want to make it a self-serve so the teams can onboard themselves, how well you have your self-serve playbooks written down, documentation on using those.
@@
 Lack of Verification of Signatures: Lack of verification of signatures is also quoted by several subjects as a main challenge.
-This arises from the factors like signing serves as a “checkmark” security tool (S4: “most people who are strictly signing all commits don’t have a meaningful process for verifying those signatures or for enforcing any policy around them.
-They just do it because”), and lack of customer interest (S14: “We have tried to … get potential customers and customers interested…but as far as I know, none of our customers…verify our signatures on images”).
+This arises from the factors like signing serves as a “checkmark” security tool (S4 (open source tooling, cloud dev tool): “most people who are strictly signing all commits don’t have a meaningful process for verifying those signatures or for enforcing any policy around them.
+They just do it because”), and lack of customer interest (S14 (POC Software & Security tool, cloud security): “We have tried to … get potential customers and customers interested…but as far as I know, none of our customers…verify our signatures on images.”)
 Verification of Artifact Before Signing: Verification generally was highlighted as a problem, especially verifying inputs to the build process of the software delivery process.
-The problem noted here is that the content of software artifacts is not readily verified before signing: as S16 put it, “if you start with the component, the component gets put into a subassembly.
+The problem noted here is that the content of software artifacts is not readily verified before signing: as S16 (security tooling, Digital technology) put it, “if you start with the component, the component gets put into a subassembly.
 The subassembly gets put into a package.
@@
 But do you know that the sum of all the pieces are accurate?
-… what are you really signing, and at what level?”.
-S11 further noted that different teams could sign arbitrary artifacts due to shared signing infrastructure.
+… what are you really signing, and at what level?”
 5.2.2 Organizational Challenges
@@
 Various teams are largely responsible for crafting their own solutions.
-As S6 said, “I can’t be more specific, but today, I’ve talked to three different teams with three different processes”.
-S15 suggested that automation could improve uniformity: “We try to break down the barriers with really good examples, by having common configuration of CI through similar GitHub workflows, and stuff like that.
+As S6 (security tooling, Digital technology) said, “I can’t be more specific, but today, I’ve talked to three different teams with three different processes”.
+S15 (internal security tooling, Aerospace Security) suggested that automation could improve uniformity: “We try to break down the barriers with really good examples, by having common configuration of CI through similar GitHub workflows, and stuff like that.
 So it’s easier for people to use.”
@@
 Subjects discussed this in terms of finances, time, and engineering effort.
-S14 said, “We need to do a lot of signing…hundreds of container images [daily]…It was a pretty significant engineering effort to build a pipeline that can handle all those signatures.”
+S14 (POC Software & Security tool, cloud security) said, “We need to do a lot of signing…hundreds of container images [daily]…It was a pretty significant engineering effort to build a pipeline that can handle all those signatures.”
 Creating Effective Signing Policy: Policies surrounding the implementation of software signing by teams are generally not considered highly effective.
 These policies include elements such as role-based access, key usage, and key issuance.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “… the other thing is policies, creating a policy that is, for one, it’s hard to do.
-And then the other part is creating an effective policy and knowing that it’s effective”
+{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3 (SAAS, SSC Security tool): “… the other thing is policies, creating a policy that is, for one, it’s hard to do.
+And then the other part is creating an effective policy and knowing that it’s effective.”
 A suggested solution to address this issue is the use of regular feedback mechanisms to assess the security policies and processes of different teams within the organization.
-As S11 said, “We want to be more proactive…internal audits, and an assessment methodology like the NIST, I think CISF.”
+As S11 (Firmware & testing software, Social technology) said, “We want to be more proactive…internal audits, and an assessment methodology like the NIST, I think CISF (NIST’s CSF).”
 No Management Incentive to Sign: A lack of management incentives was also a challenge.
-S11 said, “The investment, the time that the engineer is putting into…signing…before a release, does it get the right attention from the senior management?
+S11 (Firmware & testing software, Social technology) said, “The investment, the time that the engineer is putting into…signing…before a release, does it get the right attention from the senior management?
 Do they view it as a time well spent for the product?”
@@
 Expertise: Some subjects noted that setting up signing tools depends on the engineer’s expertise.
-Unfortunately, subjects felt that most engineers lack the necessary experience and skills for this task.
-As S9 said, “there’s a lift there as far as…expertise”.
-Developer Attitude to Signing.
-Software developers’ attitudes toward software signing pose challenges to the signing process.
+Subjects felt that most engineers lack the necessary expertise.
+As S9 (SAAS, SSC security) said, “there’s a lift there as far as…expertise”.
+Developer Attitude to Signing: Software developers’ attitudes toward software signing pose challenges to the signing process.
 Disinterest and resistance to adopting new tools are the primary attitude-related obstacles cited.
-S10 observed: “Most developers don’t really understand this stuff and don’t really want to think about it…[they are] happy if they…see a big green tick[mark]”.
-S6 added: “They refuse to get off their old tools”
+S10 (SAAS, Dev tools) observed: “Most developers don’t really understand this stuff [or] want to think about it…[they are] happy if they…see a big green tick[mark].” S6 (security tooling, Digital technology) added: “They refuse to get off their old tools.”
 Lack of Demand from Customers: Most customer-driven organizations report that software signing is challenging to implement because customers rarely request this feature in their requirements.
-S14 put it like this: “Sign[ing] our images was there from day…we haven’t found many customers that are too excited about the signing aspect”.
+S14 (POC Software & Security tool, cloud security) put it like this: “Sign[ing] our images was there from day…we haven’t found many customers that are too excited about the signing aspect.”
 5.3 RQ3: Perceived Importance of Software Signing in Mitigating Risks
-Finding 3. Subjects disagreed on this point. Some view signing as crucial, while others treat it as a secondary consideration in their overall strategy.
-Here, we aimed to understand how important practitioners perceived software signing to be in their software delivery security.
-Opinions vary on the significance of signing in the software delivery process.
-5.3.1 As Crucial to Provenance and Integrity
+In this section, we aimed to understand how practitioners perceived the importance of software signing in their software delivery security.
+Opinions on the significance of signing in the software delivery process varied.
+Subjects identified three distinct perspectives on the importance of software signing, reflecting on their personal views and how signing was treated in their current and previous roles.
+5.3.1 Signing as Crucial to Provenance and Integrity
 Many of our subjects considered provenance an integral part of security, and felt that software signing is (or will be) a guarantee of that.
+10 subjects (representing 9 organizations) had the opinion that Software signing on its own is strong enough to guarantee provenance.
 S4 said, “I think signing is massively important.
 Signing for open source is probably the single most important thing for software supply chain security”.
-5.3.2 As a Secondary Security technique
+S16’s response mirrors this, “I think [Signing’s importance] is very high…There’s a senior executive VP who has to attest that the software we release is what we say it is … he could be legally sued if the software we deliver… I mean personally, not just our organization.
+Him, because of his level in the organization, if we publish software that we have incorrectly attested to, he can be … So it’s a very high priority.”
+5.3.2 Signing as a Secondary Security Technique
 Some subjects believe that software signing is a supplementary method used to ensure the integrity of primary security techniques.
+7 subjects (6 organizations) had the opinion that Software signing does not provide enough guarantee for provenance.
 They often deemed it important only in conjunction with methods like Attestations and other Manifest forms such as the SBOM.
-As S2 said, “I think the most important thing is accurate metadata collection and centralization.
+For these organizations, signing is a valuable part of a broader, multi-layered security strategy.
+Some subjects who view signing as a strong guarantee of provenance also emphasized that adding a manifest or attestation further strengthens this guarantee.
+As S2 (SAAS, SSC Security tool) said, “I think the most important thing is accurate metadata collection and centralization.
 And even if that data’s not signed, we can still get a lot of usefulness out of it.
-Signatures, in my opinion, they offer an additional layer of security on top of what we should be doing, which is metadata collection”.
-S17 similarly views signing as a final check on top of other security practices: “Signing was essentially the way of saying, ‘I have checked, we did all the [security practices]’.”.
-5.4 RQ4: Impact of Supply Chain Security Awareness on Software Signing Practices
-Finding 4.
-1.
-Software supply chain failures usually require immediate fixes to the malfunctioning or compromised components.
-Large-scale failures typically trigger a comprehensive reassessment, leading to changes in the security strategy of a team or organization.
-2.
-Security regulations, frameworks, and standards drive the adoption of signing and other security techniques, but some subjects criticized these obligations.
-In this section, we explore how experience with software supply chain security influences a company’s decision to adopt and implement software signing.
-We asked subjects about their personal and organizational experiences with software supply chain failures (attacks, vulnerabilities, and other incidents) and how these experiences affected their team’s or organization’s software supply chain security decisions, particularly regarding software signing.
+Signatures, in my opinion, they offer an additional layer of security on top of what we should be doing, which is metadata collection.” S17 (SAAS, SSC Security tool) similarly views signing as a final check on top of other security practices: “Signing was essentially the way of saying, ‘I have checked, we did all the [security practices]’.”
+5.3.3 As a Requirement-Driven Practice
+Subjects in this category indicated that software signing was primarily used as a "check the box" technique to satisfy regulatory, framework, customer or organizational requirements.
+They did not view signing as a primary security measure.
+Instead, it was employed mainly to meet regulatory or contractual obligations, rather than being treated as an essential part of their security strategy.
+For these subjects, signing was often seen as a procedural necessity—for instance, to submit a pull request, initiate a build, or make a commit—rather than as a measure to enhance security.
+4 subjects representing 3 organizations fall into this category.
+{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S17 (SAAS, SSC Security tool): “… at certain organizations it felt like it (software signing) was more of a check-the-box thing of like, "Hey, we should sign it because that’s what people tell us to do” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S8 (Internal security tool & cloud APIs, Internet services): “…for signed binaries and code signing …a lot of those [requirements] are driven…by platform requirements…to prevent us from being blocked from…these platforms, code signing…was largely driven by [these platforms].”
+5.4 RQ4: Impact of Internal & External Events on Software Signing Practices
+In this section, we explore the impact of two kinds of security events: security failures such as Log4j, and the advent of security standards and regulations:
+  1. 1.
+    Experienced Software Supply Chain Failures: We consider these internal if a subject’s organization is a direct victim of a Software Supply Chain attack, and external if a subject’s organization is not directly impacted in a Software Supply Chain event, e.g., a general Software Supply Chain incident like the Solarwinds attack.
+  2. 2.
+    Software Supply Chain Standards: These are industry wide external events e.g., the publishing of the EO-14028 executive order on Software Supply Chain security.
+We asked subjects about their organizational experiences with software supply chain failures (attacks, vulnerabilities, and other incidents) and how these experiences affected their team’s or organization’s software supply chain security decisions, particularly regarding software signing.
 Additionally, we queried subjects on the impact of software supply chain security frameworks, standards, and regulations on their use of software signing and other security techniques.
-5.4.1 How experiences with SSC attacks influence the use of software signing
-TABLE V: Experienced software supply chain failures and type of Security change triggered.
-Type # Subjects (# Orgs) # Direct fix # $\Delta$Signing # $\Delta$General Security process
-Attack 4 (4) 4 2
-Vulnerability 9 (6) 9 1 2
-Other Incidents 5 (5) 5 2
-Software supply chain security failures typically have direct fixes, except in cases of severe attacks.
-While only 4 of our subjects indicated experiencing software supply chain attacks, only one of these attacks (where the subject was hired to help fix the aftermath) was highlighted as major, resulting in collateral damage and leading to a complete overhaul of the software and security processes.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S4: “I would say that at this point it’s probably difficult to find any process at that organization that wasn’t inspected and potentially changed as a result of that.
-Every aspect of how everything was built there was changed … There was no choice in that situation.”
-Some subjects report that software supply chain security failures led to a broader change in their team’s or organization’s security posture.
-For example, S18 describes their team’s response to the Log4j vulnerability, which was present in their system.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… it was a reality check to know that even companies like ours … can be attacked through open source libraries, … Asset inventory is the biggest thing that is lacking where you don’t know what is currently running in your environment.
-And if it is running, you don’t know how old that is, or what are the dependencies which is currently running in production.
-So that pushed us towards making sure that we have a continuous inventory of container images, which are currently running, for example, for Kubernetes, and then there are initiatives like GUAC.”
-Generally, most software supply chain reported by our subjects were not severe, and as such most of them received direct fixes.
-This could be swapping out a defaulting tool or dependency, re-releasing an affected release, publishing security advisories, etc.
-We highlight some of those responses.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S5: “So what happens is that, if the current version where that security risk is present, if that is also publicly available, we have to create a security advisory to let everyone know that if you are working with this version of our software, you should upgrade to the next version which contains the fix.
-And we let them know what the fix is and if they are affected by the fix.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “to my knowledge it did not cause a major restructuring or change that is … Yeah, I would say that they probably patched any of the security issues that they had.”
-The only instance where a signing-related fix directly resulted from a software supply chain security failure was when the subject team’s product was a safety-critical application.
-This team added intoto’s signed attestation capability.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S2: “we started reaching for the In-toto project, and we saw that it allowed you to decouple the security metadata collection and validation.
-You can abstract that away from the actual build process itself.
-So that decoupling allowed us to solve this problem of how do we know that this opaque process for the software, that opaque certification, how do we know that was followed without actually knowing the details of what that process is?”
-5.4.2 Influence of security regulations, frameworks, and standards
-TABLE VI: Regulatory and Standard Influences on Security Technique Adoption: Analysis of Subjects’ Responses. 4 Subjects say they are not influenced by either.
-ANY INFLUENCE?
-#Subjects (Yes) #Subjects (None)
-Regulation(EO-14028) 11 7
-Other Standards 8 10
-The adoption of software signing and other security techniques is heavily influenced by regulations and other standards.
-Specifically, we asked our subjects, “Are these strategies (software signing implementation & other complementary security methods(earlier listed by subjects)) influenced by regulations and standards?”.
-While answering generally, most of our subjects agreed that regulations, particularly the E0-14028, played a key role in the way their team and organization have implemented software supply chain security.
-Although software signing is not explicitly mentioned, Some subjects did note that their use of SBOMs was directly a result of regulations or standards.
-We only had 1 subject who noted that the standards influenced their use of signing in certifying attestations(SBOMs).
-S1 highlights this influence of the executive order: “Yeah, I would say certainly the collection of SBOMs and doing that more broadly has been a response to things like the executive order, at least for our team specifically.
-I would say, yeah, the executive order played just a strong role in what we are starting to do now.
-Whether or not some teams are now trying to be preventative, that’s a good question.” S7 also adds to S1’s point: “we keep the executive orders in mind when we build our product…”
-Some of our subjects (3) only relied on standards (not on regulations) for security recommendations.
-For example: {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… as an organization, currently we don’t have any government contracts.
-And I think the executive order is mostly targeted towards companies, organizations who are having contracts with the government and they need to produce certain SBOMs and things like those.
-So our push towards supply chain or better security was not influenced by the executive order.”
-S18 also adds: “But I think in terms of where we are at, I think in terms of frameworks, as you mentioned, we do look at some of these frameworks like SLSA,… There are quite a few …, and each of them target different parts of the supply chain.
-But as an organization, it’s really difficult to keep track and follow each of these standards of frameworks, to be honest, because we have limited resources, we can’t just keep on handling those.
-So we are looking at SLSA, or at least trying to get closer to where we should be for the SLSA level.
-Again, it’s not written in stone that we need to have level one, level two, level three, things like that, but that’s something which we strive to achieve sooner.”
-S14 adds to this line of thought: “The recent interests from the US federal government and software supply chain security, SBOMs, and the executive order from 2021, it’s hard to detect very much influence on our security because of those … I think when observers of software security read that executive order and then commentary in the aftermath, they often think that SBOMs have become mandatory…”
-5.4.3 Criticism of Regulations and Standards
-Also, our subjects identified several issues with the current set of software supply chain security standards and regulations that hinder their influence on the adoption of software signing and other security tools.
-We highlight four kinds of issues.
-(a) Regulations don’t specify standard data exchange format for security tools: Criticism of standards and regulations here lies in their inadequacies to specify or promote standard data interchange formats.
-This is highlighted by the lack of data formats for certain current security techniques such as Software Bill of Materials (SBOMs), other manifest-type techniques, signing tools, etc.
-The absence of standards for these techniques leads to numerous undesired practices, such as the use of vulnerability scan reports as an input for SBOM data.
-This ultimately leads to compatibility problems among different security techniques.
-{quoting}
-[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S16: “…one of the gaps with the executive order is, yeah, you give an SBOM now.
-But what are they going to do with it?
-Nobody knows yet.
-And how do you compare it?
-Because even right now, going back to the tools, there are different versions of SPDX.
-There are different versions of Cyclone DX.
-They’re the two big ones.
-This goes back to tools.
-There are a bunch of tools out there that will convert say 2.0 back to three and you come up with different answers using different tools.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “SBOMs are a good example of something that has a standard, but all of the implementations end up being different enough that they’re very difficult to translate from one format to the other or even get consistency within the same format type.”
-(b) Regulation and standards are just checkbox rules:
-Here, subjects argue that standards and regulations function more as compliance checks rather than as security enhancers.
-They point out that these standards and regulations lack security enhancement features, such as implementation models.
-Instead, they primarily promote compliance-type features, such as guidelines and best practices lists.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S17: “So there is currently a disconnect between regulations and security.
-So compliance and regulations keep people out of jail, keep people from getting fined, [while] security stops you from getting hacked.
-And the problem there is sometimes those two things can conflict.
-And compliance is there to make sure that everybody is following reasonable sets of rules and security is there to actually implement that and create the software and the processes that hit that … Regulations help provide the requirement.
-They don’t necessarily make security better on their own.
-And there’s always going to be a balance because what compliance tells you is compliance helps you with making sure that everybody’s following the standard … the only way for me to be compliant to this is to ignore the actual problem and just check the box.”
-(c) Difficult to implement in the SDLC: A group of participants highlighted the challenges encountered when attempting to integrate the recommendations of security standards into their software engineering processes.
-These challenges primarily arise from the difficulties in training engineers on the application of these standards, and the deficiencies in the technical implementation of the standards themselves.
-S12 makes this point clear in their response; {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S12: “… just having the executive order was just another way to explain it to the higher-ups… I don’t think that the executive order in particular affected the business … There’s kind of milestones.
-You don’t need to know exactly what the ins and outs are technically of achieving them … But it’s a lot of work getting that in place and upskilling all of your existing developers, et cetera.
-So yeah, anything that requires going to all 4,000, 10,000 of your developers in some of these massive organizations and making sure every single one of them does something different, it is a lot of work.”
-(d) Standards are not formalized: Criticism regarding the formalization of standards, as expressed by some participants, reflects the lack of consensus surrounding these standards.
-Participants argue that since these standards are not necessarily created by a majority of industry consortia, they should not be referred to as ‘standards’.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S13: “the term standards is imprecise.
-So some people think of standards as being very formalized but some people would say any documented practice is a standard.
-So there’s the SLSA project in it, and one might argue that that is a standard or one might argue that it’s not a standard, but it is a documented set of recommendations.
-And if people are implementing those, you could argue that they’re implementing a standard.”
+In their responses, subjects described a security failure they had experienced.
+We categorized each failure as a Vulnerability (unexploited security flaws), an Incident (SSC failures with no known malicious intent), or an Attack (deliberate and malicious failures).
+Then, subjects described changes these failures prompted in the organization.
+We categorized these as Direct Fixes (patched but no change to organizational security approach), General Security Process (large-scale changes involving multiple security techniques), and Signing (adoption or change in the implementation of software signing).
+5.4.1 Impact of Experiences with SSC Failures
+Table 5: Kinds of software supply chain failures experienced, and associated security changes.
+Type # Subjects (# Orgs) # Direct fix # Change in General Security process # Change in Signing process
+Vulnerability 9 (6) 9 2 1
+Incident 5 (5) 5 2 0
+Attack 4 (4) 4 2 0
+Table 5 summarizes these events and the resulting changes.
+Most cases were non-malicious, either Vulnerabilities or Incidents.
+However, four subjects had experienced attacks, with one major attack leading to extensive collateral damage and a complete overhaul of software and security processes.
+Across these events, the typical resolution was via Direct Fix, without an infrastructural change in the general security process or the signing process specifically.
+When they occurred, changes to the general security process typically involved adoption of a group of tooling such as for SBOMs.
+As a result of one event, a vulnerability led to changes in the signing process.
+In this case, the team (which had a safety-critical application) integrated in-toto’s signed attestation capability.
+The subject S2 (SAAS, SSC Security tool) stated: “We saw that [in-toto] allowed you to decouple the security metadata collection and validation…that allowed us to solve this problem of how do we know that this opaque process for the software, that opaque certification…was followed without…knowing the details of…that process.”
+5.4.2 Impact of Security Regulations, Frameworks, and Standards
+Our subjects reported that security frameworks, standards, and regulations had no direct impact on their organizations’ adoption of software signing as a security method.
+However, these standards and regulations did influence the general adoption of other security techniques, such as SBOM.
+9 participants from 5 organizations shared this view, while only 2 (from 2 organizations) from this group explicitly linked this to signing.
+5 participants (from 4 organizations) provided unclear statements (personal opinions or unfamiliarity with how decisions are made).
+We asked our subjects, “Are [your SSC security] strategies influenced by regulations and standards?”.
+S1 (POC software, Digital technology)’s response was typical: “Yeah, I would say certainly the collection of SBOMs and doing that more broadly has been a response to things like the executive order, at least for our team specifically…The executive order played just a strong role in what we are starting to do now.
+Whether some teams are now trying to be preventative, that’s a good question.”
 6 Discussion
+6.1 Impact of Organizational Size and Type of Produced Software on Signing Practices
+We did not observe any significant effect of organizational size (and nature of software product) on where signing was implemented (RQ1,  Figure 3).
+However, we did note patterns in the challenges reported (RQ2) and the influence of standards and regulations on the adoption of software signing (RQ4).
+It is important to note that these observations represent correlations rather than causations.
+While our findings highlight potential relationships, further study is necessary to establish causative factors and better understand the mechanisms driving these patterns.
+Effect on Challenges.
+We observed a noticeable difference in the organization size of participants who reported certain type of challenges.
+Only subjects from large organizations (S6, S8, S13, S15) talked about difficulties in operationalizing the signing process.
+We conjecture this is because of the greater coordination effort required to harmonize the engineering process of several teams.
+The nature of the software product also seem to exert some influence on the challenge experienced.
+Unsurprisingly, only subjects from organizations with non-security-focused product areas (S11 (Firmware & testing software, Social technology) and S13 (Internal security tool, Telecommunication)) reported a lack of management incentive as a barrier to signing.
+While Only participants from security-focused product teams (S3 (SAAS, SSC Security tool), S10 (SAAS, Dev tools)) reported challenges in creating effective signing policies.
+Effect on Influence of Standards and Regulations.
+As we reported, nine participants from five organizations noted that standards and regulations influenced the adoption of tools like SBOM.
+Within this group, only two participants (from two of the five organizations) explicitly linked this to signing.
+Notably, all five organizations were either large or focused on security-related products.
+6.2 Relation between Participants’ perceived importance of Signing and Signing Implementation
+We remark on a recurring dissonance between the perceived importance of Software signing by the participants and their implementation practices.
+Although all organizations signed their final products, signatures were rarely used to verify integrity or provenance Table 3.
+Ten subjects (from nine organizations) regarded Software signing as a very strong guarantee of provenance, but only three participants (from two organizations) reported verifying signatures for internal contributors, and none did so for external contributors.
+This suggests a unidirectional use of Software signing, focusing on establishing trust for their outputs while placing lower expectations on the provenance of their inputs.
+As a result, these organizations are effectively publishing software with a higher degree of provenance than they demand from their dependencies or external contributions.
 6.3 Recommendations from this study
 Improving Signature & Artifacts Verification in Signing Workflows.
-Our refined software supply chain factory model analysis has highlighted a significant gap in the current practice of signing - signature verification.
+Our refined software supply chain factory model has highlighted a significant gap in the current practice of signature verification.
 This issue was underscored by several subjects in our study.
@@
 Thus, this checkbox provides very little security properties, and requires effort to provide meaningful security gains.
-Further, while artifact verification, in this context, refers to conducting a thorough check of the integrity of the software artifact by means of identifying the signer.
-However, the use of signed attestations (i.e., are statements about a particular action or property in the supply chain) can further verify the authenticity and integrity of the process and will become crucial to further improve the security posture of organizations.
+Effective artifact verification is essential to ensure the integrity of the software.
+In this context, it involves a thorough check of the integrity of the artifact by identifying the signer.
+Moreover, the use of signed attestations—statements verifying specific actions or properties within the supply chain—can further authenticate and safeguard the process.
+Adopting these practices will be crucial for enhancing the security posture of organizations.
 Future works in this area should focus on proposing improved tooling around existing or new software signing tools to include these capabilities in their design.
 This enhancement would not only improve transparency but also reduce the costs and expertise requirements for setting up such infrastructure.
-By addressing these issues, we can ensure that signing is not just a procedural element, and instead plays a crucial role in maintaining the integrity and authenticity of the software development process.
+Also, while much of our recommendations focus on “improving tooling”, further research is needed to define what kinds of tooling would qualify.
+This research might involve experimental and empirical analyses of the current generation of software signing tools, particularly in terms of functionality, integration [65], and usability.
+As additional takeaways, we emphasize the need for further research to address critical trust gaps within the open-source ecosystem.
+Our findings ( section 5.1.4) highlight an underlying lack of trust in open-source packages, suggesting that cross-ecosystem provenance maintenance could serve as a potential solution.
+By addressing these issues, we can ensure that signing is not just a procedural element and instead plays a crucial role in maintaining the integrity and authenticity of the software development process.
 Addressing Incentive Challenges in Software Signing.
@@
 These incentives can range from support from management, the ease of using signing tools, as well as comprehensive education on the importance and application of signing.
-While some of these issues, such as the need for educating practitioners on the significance and application of signing, can be addressed with relative ease, others necessitate more elaborate planning.
+Although some of these issues, such as the need for educating practitioners on the significance and application of signing, can be addressed with relative ease, others necessitate more elaborate planning.
 For instance, fostering management support might require demonstrating the tangible benefits of software signing in terms of enhanced security and reduced risk of software tampering.
-Moreover, though the user-friendlness of signing software has been evidenced by various “why Johnny can’t” papers, more research is needed in lowering the adoption bar for software signing systems.
+Moreover, though the user-friendliness of signing software has been evidenced by various “Why Johnny Can’t” papers, more research is needed in lowering the adoption bar for software signing systems.
 Regardless of the complexity of these challenges, investing in these areas is paramount for enhancing an organization’s security posture.
@@
 This could provide valuable insights for organizations striving to improve their software signing practices.
-Criticism of Current Standards and Regulations.
-A significant finding of our study is the criticism directed towards current standards and regulations, particularly concerning their creation process and the difficulty of their implementation.
-These issues could be mitigated by ensuring that these standards and regulations are informed by a broader section of the industry consortium.
-This would involve gathering input from a diverse range of practitioners, including those from different sectors, roles, and levels of expertise.
-While frameworks like SLSA already involve collaboration among multiple organizations, we advocate for a more inclusive effort.
-This would target a larger group of practitioners, ensuring that the standards and regulations are not only comprehensive but also practical and implementable across various contexts.
-In the future, research could focus on exploring effective strategies for fostering such inclusive collaborations and assessing their impact on the quality and applicability of the resulting standards and regulations.
-By addressing these issues, we can work towards standards and regulations that are not only robust and comprehensive but also widely accepted and easily implementable.
-This would significantly enhance the security and integrity of software signing practices across the industry.
 6.4 Contrast to prior work
@@
 Distinctions of Refined Model Compared to Existing Frameworks.
-While all software supply chain security frameworks agree on the defined threat model, our refinement of this model differs from existing frameworks in several ways.
+While all software supply chain security frameworks agree on the defined threat model, our refinement of this model differs from existing ones in several ways.
   1. 1.
@@
 7 Conclusion
-Signature-based supply chain provenance is a topic of increasing importance, driven by cyberattacks, executive orders, and new industry standards.
-Although the literature contains many studies of signing as a technology, there have been few studies of the use of signing in engineering contexts.
-Those that exist have focused on open-source, or on the mechanics of signing.
-In this paper, we present the first qualitative examination of signing in engineering organizations’ security strategies.
-We interviewed 18 high-ranking practitioners across 13 organizations, including C-suite staff, principal and senior engineers, managers, and consultants.
-Our four main findings are that (1) Software products are being signed, but the signatures are not yet being validated; (2) The factors that hamper signing include infrastructure and compatibility issues, lack of expertise, developer and organizational attitudes, and lack of resources; (3) Signing is considered an important guarantee of provenance, but it is not the main technique these organizations consider in their cybersecurity postures; and (4) Organizational behaviors are being influenced by major failures (e.g., SolarWinds) and by regulations and industry standards, but some of our subjects expressed concern about the development process and quality of those recommendations.
-Our results indicate many opportunities for technical and operational advances in the use of software signing to improve software supply chain provenance.
-8 Replicability
-For anonymity reasons, we do not plan to share the raw transcripts of our interviews. However, our interview protocol, codebook, etc. are available in the Appendices.
+Software signing is widely recommended for ensuring software provenance, yet many organizations face challenges when integrating it into their processes.
+In this first qualitative study, we examine the practices and challenges of software signing through interviews with 18 senior practitioners across 13 organizations.
+We identified that while software products are signed, the signatures often go unvalidated.
+Key barriers include infrastructure limitations, lack of expertise, and insufficient resources.
+Although signing offers strong provenance guarantees, it remains secondary in many organizations’ cybersecurity strategies.
+Additionally, organizational behaviors are influenced by major incidents like SolarWinds and industry regulations and standards, with concerns raised about the quality and development of these standards.
+Our findings shed light on the real-world practices, challenges, and organizational factors shaping software signing adoption, offering guidance for organizations on their journey toward more secure software supply chains.
 8 Acknowledgments
-We thank the study participants for contributing their time to this work. Taylor Schorlemmer and Wenxin Jiang assisted in typesetting. We acknowledge support from Google, Cisco, and NSF #2229703.
-Appendix A Outline of Appendices
-The appendix contains the following material:
-  • Appendix B: The interview protocol.
-  • Appendix C: The codebooks used in our analysis, with illustrative quotes mapped to each code.
-  • Appendix D: Analysis of saturation in our data.
-  • Appendix E: Uncommon challenges, omitted from the main paper due to space limitations.
-Appendix B Interview Protocol
-Table I gave a summary of the interview protocol.
+We thank the study participants for contributing their time to this work.
+Taylor Schorlemmer, Sofia Okorafor, and Wenxin Jiang assisted in typesetting.
+We acknowledge support from Google, Cisco, and NSF #2229703.
+9 Ethical Considerations
+This work considers the human and organizational factors governing the adoption of software signing techniques in software engineering work.
+To elicit this data, we developed a research protocol, including subject recruitment, interview content, and data processing and analysis.
+This protocol was approved by our organization’s Institutional Review Board (IRB) prior to the study conduct.
+All members of the research team who interacted with subjects or their non-anonymized data have completed our organization’s training procedures for handling personally identifiable information.
+Data from individual subjects were anonymized following best practices for qualitative data analysis.
+Any human-subjects research effort must weigh the risks and benefits for individual research subjects as well as for society.
+The primary risk for subjects is the hazard to their individual or organizational reputation when describing what might be interpreted as sub-standard cybersecurity practices.
+We therefore anonymized subjects by rank and by organization.
+A secondary cost for subjects was their time.
+Subjects were not coerced to participate, and we offered an incentive of \$100 for their participation.
+The primary potential benefit for our subjects, their organizations, and the societies in which they operate, is that the resulting data may motivate improvements in the software signing ecosystem, as we discussed in section 6.
+In our judgment, the potential benefit to software cybersecurity far outweighs the risks posed to our subjects and their organizations.
+We therefore attest that:
+  • The research team considered the ethics of this research, that the authors believe the research was done ethically, and that the team’s next-step plans (e.g., after publication) are ethical.
+10 Compliance with Open Science Policy
+We acknowledge the commitment to open science, where authors are generally expected to share their research artifacts openly. The research artifacts associated with this study are
+  • Raw transcripts of interviews
+  • Anonymized transcripts of interviews
+  • Interview protocol
+  • Codebook
+Things we have shared: The interview protocol is a crucial part of any interview study, since it allows for the critical review of a study design as well as its replication.
+Our full interview protocol is included in Appendix E.
+Since it is semi-structured, we include all of the questions asked of all subjects, as well as examples of the follow-up questions we asked.
+We also share the codebook, with codes, definitions, and example quotes that we coded for each code — see Appendix F.
+Our full codebook is available at https://doi.org/10.5281/zenodo.14660193.
+Things we cannot share: For subject privacy reasons, and for IRB compliance, we cannot share the raw transcripts.
+We are also unwilling to share the anonymized transcripts of the interviews.
+Given the high organizational ranks of many of our subjects, and the small size of the subject pool resulting from our recruiting strategy, we believe there is a high risk of de-anonymization even of anonymized transcripts.
+Therefore, we cannot share the anonymized transcripts.
+Appendix A Traditional Software Signing Workflow
+We show a typical traditional package signing workflow in Figure 4.
+Figure 4: Typical workflow for software signing and verifying signatures.
+The component author packages (A) and signs (B) their software.
+The signed package (C) and public key (D) are published.
+To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
+Appendix B Organizational Demographics
+We summarize additional organizational demographic information of our subjects in Table 6.
+This table provides further context for the subject’s organizational size, type of software product, and how many of our subjects belonged to each organization.
+Table 6: Organizational Demographics. To enhance anonymity, We only highlight the number of subjects in each organizational category. letters A-L is used to depict each organization.
+Type Breakdown (#Organizations|#Subjects)
+Organizational Size (Employee Size) Small (<100) (4/6), Medium (<1500) (3/4), Large (>1500) (6/8)
+Product Area Digital technology (1/3), SSC Security (2/4), Social technology (1/1), Dev tools (1/1), Telecommunications (1/1), Cloud security (2/3), Aerospace Security (1/1), Internet services (2/2), Cloud + OSS Security (1/1), Cloud + Dev toools (1/1)
+Subject Distribution A (3), B (3), C (1), D (1), E (2), F (1), G (1), H (1), I (1), J (1), J (1), K (1), L (1)
+Appendix C Code Saturation Curve
+We present the trends from the data analysis of our interview subjects in Figure 5.
+The total number of codes per interview is shown by the blue trend line, with a median of 84.5 codes per interview.
+Participant S8 had the fewest codes (50), while S7 had the most (130).
+The orange trend line represents the unique number of codes in each interview, with a median of 33 unique codes. Subject S4 had the fewest unique codes (24).
+The green trend line illustrates the saturation curve for the study, indicating that no new unique codes were observed beyond interview 11 (S11).
+As noted earlier (section 4.3), the pilot interviews (S1 and S2) were included due to minimal changes in the interview protocol.
+These interviews also contained more unique and total codes than the study’s median values, further justifying their inclusion.
+Figure 5: Saturation curve.
+Interviews are plotted in the order in which they were conducted.
+Each interview covered many topics in detail (orange line, blue line).
+However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
+Appendix D Other Less Reported Technical Challenges
+D.0.1 Managing Signed Artifacts
+Managing signed software artifacts post-signing should follow good practices, but this is currently a challenge.
+{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S9: “Also managing all the artifacts that are signed … It’s great that you sign things, but what do you do with it afterwards?
+I think that’s kind of something that we’re facing right now, especially when it comes to container management.
+Great, we signed containers, but now what?
+… Okay, great, now what?
+All those downstream things that we’re trying to figure out.”
+D.0.2 Identity Management/Authentication
+Managing and authenticating the identities of team members authorized to sign at various stages of the software delivery process poses challenges in some instances.
+{quoting}
+[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “in general one of the challenges that I hear from other security architects and security leads, and program managers, and people at that level, is key management and identity management as being a barrier for signing and security and authentication in general.”
+Possible solutions suggested and used by some participants include the use of short-lived or constant key rotation.
+{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S7: “We really rely on short-lived tokens for authentication.
+I think that is one way that we really try to restrict access to our system and make sure that there’s no long-lived secrets, there’s no long-lived access tokens anywhere in our system.
+For any signing that we do, I think those keys are rotated quite frequently.
+So, that’s another thing we do to prevent any attack should key information be leaked.
+I’m trying to think what else.”
+Appendix E Interview Protocol
+Table 1 gave a summary of the interview protocol.
 Here we describe the full protocol.
@@
 We include Section D of the protocol for completeness, although it was not analyzed in this study.
-TABLE VII: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions.
+Table 7: Our full interview protocol.
+The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions.
+We did not analyze part D in this study.
+We excluded this data due to limited detail and confidentiality concerns.
+Following qualitative reporting best practices  [85], we disclose this topic to acknowledge its potential influence on data collection.
 Themes Sub-themes Questions
@@
 Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
-C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from  LABEL:sec:_B-sec-experienced): why did the team choose to implement software signing?
+C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from Section B): why did the team choose to implement software signing?
 Is software signing the team’s major strategy to secure its supply chain?
@@
 Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
-C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism?
+C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism? (This question is derived from  [39]).
 D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
@@
 Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
-D-6: Have you/your team considered switching this tool for another? AWhat other tools have been considered or are currently being considered?
-Appendix C Codebooks
-Table VIII describes our codebook, with code, definition, and example quote tagged with that code.
-TABLE VIII: Excerpts from the final codebook we used for our analysis. Sample codes under each theme give examples. Note that the theme, “Others and Specific tooling” is not discussed in the paper.
+D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
+Appendix F Codebooks
+Table 8 describes our codebook, with code definition, and example quote tagged with that code.
+Table 8: Excerpts from the final codebook we used for our analysis. Sample codes under each theme give examples. Note that the theme, “Others and Specific tooling” is not discussed in the paper.
 Code Definition Sample Quote
@@
 Choice of tool - General Reasons What influences choice of software supply chain security tools except signing tools S17: "A lot of the tools are not particularly easy to use "
-Appendix D Saturation Analysis
-Figure 5 shows the saturation curve associated with our interviews.
-Figure 5: Saturation curve.
-Each interview covered many topics in detail (orange line, blue line).
-However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
-Appendix E Uncommon Challenges
-E.0.1 Managing Signed Artifacts
-Managing signed software artifacts post-signing should follow good practices, but this is currently a challenge.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S9: “Also managing all the artifacts that are signed … It’s great that you sign things, but what do you do with it afterwards?
-I think that’s kind of something that we’re facing right now, especially when it comes to container management.
-Great, we signed containers, but now what?
-… Okay, great, now what?
-All those downstream things that we’re trying to figure out.”
-E.0.2 Identity Management/Authentication
-Managing and authenticating the identities of team members authorized to sign at various stages of the software delivery process poses challenges in some instances.
-{quoting}
-[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “in general one of the challenges that I hear from other security architects and security leads, and program managers, and people at that level, is key management and identity management as being a barrier for signing and security and authentication in general.”
-Possible solutions suggested and used by some participants include the use of short-lived or constant key rotation.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S7: “We really rely on short-lived tokens for authentication.
-I think that is one way that we really try to restrict access to our system and make sure that there’s no long-lived secrets, there’s no long-lived access tokens anywhere in our system.
-For any signing that we do, I think those keys are rotated quite frequently.
-So, that’s another thing we do to prevent any attack should key information be leaked.
-I’m trying to think what else.”
+Appendix G Code Evolution Diagram
+Figure 6 describes the evolution of our codebook over five rounds of revisions.
+Figure 6: Codebook Evolution: The codebook underwent five rounds of revisions, with changes highlighted by different colors: round 2 (pink), round 3 (blue), round 4 (green), and round 5 (yellow).
+Arrows indicate the merger or replacement of codes.
+The final codes from round 5 were used in this paper, and categorized thematically.
+We combined framework and thematic analyses to develop the codebook.
+First, we induced codes from each transcript and analyzed them thematically.
+During the framework analysis, we mapped codes to the four stages/blocks of the software supply chain factory model.
+Round 1: The codebook was defined based on emerging themes.
+Round 2: After the intercoder reliability assessment, two categories were merged.
+Round 3: Blue-highlighted codes (6) were added after memoing and coding the second set of transcripts.
+Round 4: Green-highlighted codes (4) were added.
+Round 5: Analysts A1 and A2 determined that all new codes from stages 3 and 4 (except three(3)) were minor and did not fit the established categories.
+We identified the following themes through framework analysis: Build, Source Code, Deployment, and Third-Party Dependency.
+Additionally, through thematic analysis, we derived themes such as the Effect of SSC Failures on Signing Adoption, Perceived Importance of Signing, Influence of SSC Standards and Regulations on Signing, and Challenges.